### v1

Abstract
We investigate the optimal charging strategy for large-scale electric vehicles in smart grids using a finite-horizon framework within a mean field game approach. By proving the existence and uniqueness of the solution to the consistency condition equation, we analyze the optimal charging strategies for electric vehicles in the mean field limit case. We then construct the approximate optimal charging strategies for a finite number of vehicles in both non-cooperative and cooperative games, and give a numerical experiment accordingly. Lastly, we present numerical analyses to illustrate approximate strategies of non-cooperate and cooperate game.
Keywords: Smart grid, mean field game, mean field equilibrium, approximate Nash equilibrium.
MSC 2020: 91A06, 49L12, 49N80, 60H30
1 Introduction
As the global energy structure continues to shift towards low-carbonization, the number of electric vehicles (EVs) is growing rapidly. However, the spatiotemporal concentration of large-scale EV charging loads presents a significant challenge to the supply-demand balance of smart grids. In this context, developing a scientifically robust and highly efficient collaborative optimization strategy framework for EV charging is not only crucial for enhancing the operational resilience of the smart grid but also a core breakthrough for optimizing users’ electricity costs and achieving efficient allocation of energy resources (Sultan et al. (2022) and Yetkin et al. (2024)). Typically, the number of EVs requiring charging in the grid is enormous, leading to the “curse of dimensionality” when determining the optimal charging (or storage) strategy for each vehicle.
In the context of large-scale network structures, mean field games (MFGs) were independently proposed by Huang et al. (2006) and Lasry and Lions (2007), which devoted to the analysis of dynamic systems where a large number of players interact strategically with each other. By employing the “mean field” approximation, MFGs simplify the complex multi-player game problem into an interaction between a single player and the aggregate behavior of the population. This framework provides an effective analytical tool for addressing the interactions among players in large-scale systems. For more comprehensive insights into MFG theory, the reader may refer to Bensoussan et al. (2016), Carmona and Delarue (2018), Gomes and Saúde (2014) and the additional references cited within those works. Its application has spanned a wide range of fields including economics and finance  (Lacker and Zariphopoulou, 2019; Nuno and Moll, 2018), large networks or graphs (Lacker and Soret, 2022), operations research (Wang and Huang, 2019; Wang et al., 2022), engineering and machine learning (Zhou and Xu (2022)), epidemics control(Laguzet and Turinici (2015); Lee et al. (2021)), smart grid (Gomes and Saúde (2021)) and more. For the tractable case, the MFG method is also intensively studied in the LQG framework, due to the elegance of its analytical tractability. In particular, Huang (2010) introduces LQG games that feature a major player with substantial influence and a multitude of minor players. Subsequently, Bensoussan et al. (2016) provide a comprehensive study of a general class of MFGs in the LQ framework. Zhang and Li (2019) explore linear-quadratic MFGs described by forward-backward stochastic differential equations. Gomes et al. (2023) study the case of linear dynamics for the supply and a cost function consisting of price, control and commodity value, which extends the LQG framework by introducing the product of control and price in the cost function. MFGs offer a natural and elegant mathematical approach to these challenges, delivering decentralized, scalable, and theoretically grounded solutions for optimizing the performance of smart grids.
Most of the MFG literature in smart grids focuses on two types of problems: Nash equilibria or social optimum. Nash equilibria in mean-field control has been investigated in Saldi et al. (2019) and Guo et al. (2022). The closest work to ours, authored by Cohen and Zell (2025), examines finite-state, infinite-horizon mean-field games with ergodic costs, demonstrates that solutions to the mean-field game system yield approximate Nash equilibria in corresponding finite-player games, and establishes a large deviation principle for empirical measures associated with these equilibria.
Social optimum solutions for MFG are also well documented. For example, Li et al. (2016) examine the relationship between mean-field games and social welfare optimization problems. Salhab et al. (2016) focus on dynamic collective choice problems, and address scenarios where a large number of players cooperatively choose between multiple destinations while influenced by group behavior. Feng et al. (2021) involve a major agent and numerous minor agents and investigates a mixed stochastic LQG social optimization. Moreover, Wang and Huang (2019) consider both the non-cooperative and cooperative solutions for a system with sticky prices and adjustment costs. It proposes an auxiliary limiting optimal control problem subject to consistent mean field approximations, and the method is enlightening to our work.
Moreover, versatile framework for modeling price dynamics in markets with numerous interacting agents are applied in MFGs. Notable approaches are listed below: Gomes and Saúde (2021) introduce a price formation model where numerous small players can store and trade a commodity like electricity; Gomes et al. (2020) propose a MFG model for price formation of a commodity with production subject to random fluctuations. The dynamic game models have been extensively utilized to analyze duopolistic competition with sticky prices. Existing literature related to it includes Cellini and Lambertini (2004), Wiszniewska-Matyszkiel et al. (2015), Valentini et al. (2021), Kańska and Wiszniewska-Matyszkiel (2022) and Hoof (2021). Besides, Cellini and Lambertini (2004) and Wiszniewska-Matyszkiel et al. (2015) propose an evolution law for the sticky price in dynamic oligopoly, which will be applied to our work.
This paper aims to develop a strategic framework for a large market composed of numerous agents, using the methodology of MFG. Within a setting characterized by sticky prices and finite time horizons, we investigate how agents can optimize their behavior to minimize expected losses. In the cooperative game case, our objective is to reduce the average social cost across the market. On the other hand, under the non-cooperative framework, we focus on constructing an approximate Nash equilibrium that captures the decentralized decision-making of individual agents. This dual perspective allows for a comprehensive analysis of collective and competitive behaviors in large-scale smart grid systems. Given the close resemblance between our model and the models of Gomes et al. (2023) and Wang and Huang (2019), we now emphasize the distinctions in both model structure and methodological approach. The model in Gomes et al. (2023) determines the mean process of the control variable through a stochastic differential equation. Subsequently, by employing variational methods, they seek a price process such that, when agents take optimal actions to minimize transaction costs, the market clears and supply meets demand. Unlike in their work, our price process is endogenously generated within the system and is a stochastic process driven by control variables. The control variables can be freely adjusted, and the agents’ commodities are subject to independent noise disturbances. The more fundamental difference lies in our objective: we aim to construct $\epsilon$-optimal solutions for both cooperative and non-cooperative games in a large system, rather than achieving supply-demand equilibrium. Besides, our methodological approach mainly involves approximating stochastic processes using fixed functions, rather than employing variational techniques. Different from Wang and Huang (2019), which aims to find an $\epsilon$-Nash equilibrium and $\epsilon$-socially optimal solution for the system based on cost minimization within an infinite-horizon framework, we consider a finite-horizon setting and adopt a cost function of the LQG form, which deduces to the feedback strategies that depend on the state process. More importantly, our cost function is positively correlated with the product of price and demand, while this term is negatively correlated with the cost function proposed by Wang and Huang (2019). This positive correlation introduces great amount of complexity to the problem, as it implies that increasing the norm of the control variable can significantly reduce one component of the cost function to a very small negative number, potentially undermining the coercivity condition. Overcoming the complications introduced by this correlation constitutes the principal difficulty of the present work and serves as the guiding objective behind many of the subsequent proofs.
Our mathematical contributions can be summarized in three aspects: First, within a finite-horizon framework, we propose a mean-field interaction cost function, through which we derive approximate solutions for both the cooperative and non-cooperative games. To the best of our knowledge, it is relatively rare in the literature to address both cooperative and non-cooperative formulations simultaneously in a finite-horizon framework. Second, for specific matrix equations, we construct the solution induced by specific initial values and study the existence and uniqueness of the mean field through its properties. Most of the existing literature employs the Banach fixed-point theorem or other methods to find equilibrium points. In contrast to these approaches, our method is able to more precisely characterize the state of the mean field. Third, we apply the Positive Real Lemma to demonstrate that certain parameter restrictions can ensure that the coercivity condition holds, thereby effectively addressing the challenge posed by the positive correlation between the cost function and the product of price and demand term. Although Wang and Huang (2019) also employs the Positive Real Lemma, the problem we consider is more complex, and the manner in which we utilize the lemma is correspondingly more intricate.
The rest of the paper is organized as follows: In Section 2, we introduce the model of the smart grid, which contains a large number of electric vehicles. We then construct the $\epsilon$-Nash equilibrium charging strategies for the EVs in the non-cooperate games. Afterwards, we construct the asymptotic social optimum charging strategies for the cooperate games. In Section 3, we give numerical simulations for the charging strategies proposed in both the non-cooperate games and cooperate games. Section 4 concludes.
2 The Model and the Optimal Strategies
In this section, we propose the sticky price model and the optimal charging strategies of smart grid with numerous agents (e.g. EVs). First, we introduce our model. Let the number of agents in the smart grid be $N$. The agents control their electricity (also referred to as commodities) through controlling the charging rate over a shared common time horizon $[0,T]$. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions, which supports the independent $N$-dimension standard Brownian motions $(W_{t}^{1},\cdots,W_{t}^{N})_{t\in[0,T]}$. Denote the electricity held by agent $i$ at time $t$ by $X_{t}^{i}$. The dynamic of $X^{i}$ satisfies
$dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i}\in\mathbb{R}.$ (2.1)
Here, the parameter $\sigma_{i}\geq 0$ measures the volatility of agent $i$’s electricity, and the stochastic noise term represents uncertainty or variability in the smart grid system, such as grid disturbances or physical noise, communication or control delays, or user behavior uncertainty. In addition, we assume that agent $i$ is equipped with an initial electricity $x_{0}^{i}\in\mathbb{R}$, As shown in (2.1), the control variable for agent $i$ is denoted by $v^{i}$, also referred to as trading rate (or charging rate).
We then introduce the evolution of the price process $P$. We consider a dynamic oligopoly setting, where prices do not instantly adjust to changes in market conditions, such as shifts in supply or demand. Instead, its evolution is affected by its current price and average trading rate. A natural way of modeling the sticky price in dynamic oligopoly is (Cellini and Lambertini (2004) and Wiszniewska-Matyszkiel et al. (2015))
$dP_{t}=\alpha(\beta-Q_{t}-P_{t})dt,\quad P_{0}=p_{0},$ (2.2)
where $\alpha,\beta>0$ are constant market factors, and the initial price is $p_{0}$. The parameter $\alpha$ measures the sensibility of $P$ to current situations. $Q_{t}$ is the average trading rate at time $t$. In this paper, it is equal to $\frac{1}{N}\sum_{i=1}^{N}v_{t}^{i}$.
In the following, by referring to existing literature, we propose a reasonable cost function for each agent in the game. Inspired by Gomes et al. (2023), we let the expected overall cost of agent $i$ be of the following form
$J_{i}(v^{i},\boldsymbol{v}^{-i})=\mathbb{E}\left[\int_{0}^{T}\left(L(X_{t}^{i},v_{t}^{i})+P_{t}v_{t}^{i}\right)dt+\Psi(X_{T}^{i})\right],$ (2.3)
where $\boldsymbol{v}^{-i}:=\{v^{1},\cdots,v^{i-1},v^{i+1},\cdots,v^{N}\}$. We assume the running cost and the terminal cost to be
$L(x,v)=\frac{\eta}{2}(x-\kappa)^{2}+\frac{c}{2}v^{2}\quad\mathrm{and}\quad\Psi(x)=\frac{\gamma}{2}(x-\zeta)^{2}$
respectively. The parameter $\zeta$ corresponds to the preferred final storage, and $\kappa$ is the preferred instantaneous storage. We make the assumption that $\eta>0,\ \gamma\geq 0,\ c>0,\ \kappa\in\mathbb{R},\ \text{and}\ \zeta\in\mathbb{R}$ are constants. To introduce the admissible strategy set of $v^{i}$, we assign every stochastic control $w$ the norm $\left\|w\right\|=\sqrt{\mathbb{E}\left[\int_{0}^{T}w_{t}^{2}dt\right]}$. Then, $v^{i}$ should belong to the admissible strategy set
[…]
Finally, building on the previously proposed cost function, we introduce an objective function for the system. We consider two different game cases. In one case (namely, the non-cooperate game case), each agent controls its commodity independently, and does not take into account how its actions might affect others’ expected cost. In other words, agent $i$ only aims to minimize its expected value function $J_{i}$, and does not consider how $J_{k}(\forall 1\leq k\leq N,k\neq i)$ changes according to its actions. Under this setting, it is natural for us to find an $\epsilon$-Nash equilibrium for the system.
A Nash equilibrium is a key concept in game theory that represents a stable state in a strategic interaction where no player can improve their outcome by unilaterally changing their strategies. It occurs when each player’s strategy is optimal given the strategies chosen by all other players. However, as $N$ enlarges, it becomes rather challenging to find the exact Nash equilibrium solution, so we aim to find a set of strategies that approaches the best outcome as $N$ enlarges, referred to as a set of $\epsilon$-Nash equilibrium strategies. An $\epsilon$-Nash equilibrium characterizes a situation where each player’s chosen strategy is almost the best response to the strategies of the others, with the difference between the expected personal cost of this strategy and that under the optimal strategy not exceeding a positive number $\epsilon_{N}$, which vanishes as $N$ goes to $\infty$. We introduce the precise definition of an $\epsilon$-Nash equilibrium:
Definition 2.1.
A set of admissive strategies $\hat{\boldsymbol{v}}=\{\hat{v}^{1},\cdots,\hat{v}^{N}\}\in\mathcal{A}^{N}$ is an $\epsilon$-Nash equilibrium if
$J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})-\epsilon_{N}\leq\inf_{v^{i}\in\mathcal{A}}J(v^{i},\hat{\boldsymbol{v}}^{-i})\leq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\quad i=1,\ldots,N,$ (2.4)
where $\epsilon_{N}$ goes to 0 as $N\to\infty$.
In order to further characterize the rate of $\epsilon$ reduction, we introduce the following notion. A number dependent on $N$ is called an $O(\frac{1}{\sqrt{N}})$ number if there exists a fixed constant $M$ (depending on all the parameter introduced, including $p_{0}$, $T$, and others), which is independent of $N$, such that for $N$ large enough (i.e., there exists a constant $N_{0}>0$, such that for $\forall N>N_{0}$), this number’s absolute value does not exceed $\frac{M}{\sqrt{N}}$. Similarly, we can define $O(1)$ numbers and $O(\frac{1}{N})$ numbers. Apart from that, when we state that a certain quantity dependent on $N$ (for example, a function or a variable $h(N)$) satisfies $h(N)\leq O(\frac{1}{\sqrt{N}})$, we mean that there exists an $O(\frac{1}{\sqrt{N}})$ number, such that $h(N)$ is always no greater than this $O(\frac{1}{\sqrt{N}})$ number. We introduce a similar form of comparing rule for $O(1)$ and $O(\frac{1}{N})$ numbers in a completely analogous way.
Under the second case (namely, the cooperate game case), instead of controlling the commodities independently, the agents cooperate with each other, and aim to achieve the best collective outcome. In this article, the collective outcome represents the average value of $J_{i}(v^{i},\boldsymbol{v}^{-i})$. It reads
$J_{soc}(v^{1},v^{2},\cdots,v^{N})=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(L(X_{t}^{i},v_{t}^{i})+P_{t}v_{t}^{i}\right)dt+\Psi(X_{T}^{i})\right].$
In large-scale systems, finding the exact social optimum becomes complex, so we focus on finding a set of strategies that approaches the best collective outcome as the number of agents tends to infinity. More specifically, we introduce the following definition:
Definition 2.2.
We say that a set of admissive strategies $\{\check{v}^{i}\}_{i=1}^{N}$ is a set of $\epsilon$-optimal social strategies if
$|J_{soc}(\check{v}^{1},\check{v}^{2},\cdots,\check{v}^{N})-\inf_{v^{i}\in\mathcal{A},1\leq i\leq N}J_{soc}(v^{1},v^{2},\cdots,v^{N})|\leq\epsilon_{N},$ (2.5)
where $\epsilon_{N}$ must vanish to 0 as $N\to\infty$.
The restriction for $\epsilon_{N}$ reflects that the expected cost under the proposed strategies converges to that of the best strategies as $N$ goes to $\infty$.
It is worth emphasizing that asymptotic social optimum strategies may not be $\epsilon$-Nash equilibrium strategies. Under the asymptotic social optimum strategies, an agent may significantly reduce its own expected cost by only changing its own strategy, while causing even more cost for others, and comprehensively, give rise to the overall expected cost.
One more point needs to be stated in advance. In this paper, when we solve first-order ordinary differential equations, unless otherwise specified, we are looking for classical solutions (i.e., continuously differentiable solutions), not weak or viscosity ones.
2.1 Non-cooperate Game
In this subsection, we aim to find an $\epsilon$-Nash equilibrium for this system. We will use a mean-field approximation approach, in which we construct $C[0,T]$ functions to approximate certain processes. Our approach can be outlined in the following steps.
  1. 1.
    The first step in constructing a mean-field approximation is to approximate $Q$ with a continuous function $\bar{Q}=(\bar{Q}(t))_{t\in[0,T]}$. It follows from (2.2) that the approximate price process $(\bar{P}(t))_{t\in[0,T]}$ satisfies
    $d\bar{P}(t)=\alpha(\beta-\bar{Q}(t)-\bar{P}(t))dt,\quad\bar{P}(0)=p_{0}.$ (2.6)
  2. 2.
    Replace $P_{t}$ by $\bar{P}(t)$ in objective function $J_{i}$ to obtain an auxiliary objective function
    $\bar{J}_{i}(v^{i};\bar{P})=\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right],$ (2.7)
    subject to state process $(X_{t})_{t\in[0,T]}$, which satisfies (2.1).
  3. 3.
    For each agent $i$, solve the auxiliary control problem (2.7) to find the optimal control $\hat{v}^{i}$. Calculate the corresponding average optimal control and remove the noise term to obtain a determinist process. By requiring that this process equal $\bar{Q}$, we determine the exact form of $\bar{Q}$.
  4. 4.
    Prove that under certain assumptions, the set of controls derived from $\bar{Q}$ is indeed an $\epsilon$-Nash equilibrium.
Remark 1.
The rationale for proposing the first two steps is primarily based on the following reasons. Firstly, when the population enlarges to $\infty$, the representative agent has no influence on $Q$, and therefore has no influence on $P$, as but one agent amid a continuum. From its perspective, the process $P$ should be evolved in its cost as a deterministic function. Secondly, the optimization problem for (2.3) becomes much easier to solve once we treat $Q$ and $P$ as fixed continuous functions, instead of random ones. For each $\bar{Q}$, a unique process $\bar{P}$ is determined by (2.6). Each agent determines its optimal control and gives the corresponding average investment process according to $\bar{P}$. If the expectation of this average investment process is equal to $\bar{Q}$, a closed-loop is formed. Within this loop, the strategy adopted by each agent yields outcomes that closely approximate those of the actual optimal strategy as the number of agents increases.
For the given process $\bar{Q}\in C[0,T]$, the approximate price process $\bar{P}$ is uniquely determined by (2.6). We propose the value function of the auxiliary control problem (2.7),
$K_{i}(t,x)=\inf_{v^{i}\in\mathcal{A}}\mathbb{E}\left[\int_{t}^{T}(L(X_{s}^{i},v_{s}^{i})+\bar{P}(s)v_{s}^{i})ds+\Psi(X_{T}^{i})\right].$
where $t$ represents the starting time, and $x$ represents the initial value of $X^{i}$. By dynamic programming principle, the HJB equation satisfies
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$ (2.8)
along with the terminal condition $K_{i}(T,x)=\frac{\gamma}{2}(x-\zeta)^{2}$. We make the ansatz ${K}_{i}(t,x)=a(t)x^{2}+B(t)x+F_{i}(t)$ yields. Furthermore, the optimal feedback control achieving the minimum is given by
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)x+B(t)}{c},\quad\forall t\in[0,T].$ (2.9)
Plug (2.9) into (2.8) to obtain the following ODE system:
$\qquad\qquad\qquad a^{\prime}(t)+\frac{\eta}{2}-\frac{(2a(t))^{2}}{2c}=0,$ $a(T)=\frac{\gamma}{2},$ (2.10)
$\qquad\qquad\qquad B^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{P}(t)+B(t))}{c}=0,$ $B(T)=-\gamma\zeta,$ (2.11)
$\qquad\qquad\qquad F^{\prime}_{i}(t)-\frac{(\bar{P}(t)+B(t))^{2}}{2c}+\frac{\eta\kappa^{2}}{2}+\sigma_{i}^{2}a(t)=0,$ $F_{i}(T)=\frac{\gamma\zeta^{2}}{2}.$ (2.12)
To make the ansatz hold, we will verify that (2.10),(2.11) and (2.12) give a solution. We thereby introduce subsection 2.1.
Proposition 2.1.
The solution for (2.10), (2.11) and (2.12) exists and is unique.
The proofs of the Propositions in this paper have been left to Appendix B. We next present the verification theorem on the best response control for minimizing (2.7), which demonstrates that our ansatz holds.
Theorem 2.2.
(Verification theorem) We introduce the feedback control function as follows:
$\hat{v}_{t}^{i}(t,X_{t}^{i})=-\frac{\bar{P}(t)+2a(t){X}_{t}^{i}+B(t)}{c},\quad\forall t\in[0,T].$
Then, the SDE
$d{X}_{t}^{i}=\hat{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad{X}_{0}^{i}=x_{0}^{i}$
admits a unique solution, denoted by $\hat{X}_{t}^{i}$, and $\hat{v}^{i}$ is an optimal Markovian control for minimizing (2.7).
Proof.
$\hat{v}_{t}^{i}$ is a measurable function w.r.t $(t,x)\in[0,T]\times\mathbb{R}$, and satisfies the terminal condition $K_{i}(T,\cdot)=\Psi(\cdot)$. From the proof of subsection 2.1, $a(t)$, $B(t)$ and $F_{i}(t)$ are bounded on $[0,T]$, so $K_{i}(t,x)$ satisfies the quadric growth condition (i.e. there exists $C>0$ such that $|K_{i}(t,x)|\leq C(1+x^{2}),\ \forall(t,x)\in[0,T]\times\mathbb{R}$). Besides, (2.10)-(2.12) imply that, (2.9) solves the equation
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}$
$=\frac{\partial{K}_{i}(t,x)}{\partial t}+\hat{v}_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({\hat{v}_{t}^{i}})^{2}+\hat{v}_{t}^{i}\bar{P}(t)+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
The result then follows from Pham (2009). ∎
Now, we will approximate the average commodity process with a deterministic process. The underlying reason for this approximation can be explained as follows: by denoting the average of $X_{t}^{i}$ by $\overline{X}_{t}$, the average of this optimal control has the form $\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c}$, which leads to the mean commodity process evolving as follows
$d\overline{X}_{t}=-\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c}dt+\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i},\quad\overline{X}_{0}=\overline{x}_{0}^{N}.$
It is worth noticing that $\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i}$ (the average of the independent Brownian motions) forms a Brownian motion with a small magnitude, denoted by $\sigma dB_{t}$, where $\sigma:=\frac{1}{N}\sqrt{\sum_{i=1}^{N}\sigma_{i}^{2}}$ represents its magnitude and $B_{t}$ can be interpreted as the standardized form of the Brownian motion obtained by aggregating the independent noise. It is easily checked that this magnitude gradually decreases to zero as $N$ increases, provided that $\{\sigma_{i}\}_{i=1}^{N}$ is bounded. Therefore, by using a mean-field approximation, among which we let $N\to\infty$, we will approximate the average commodity process with a determinist function $\bar{x}(t)$, whose dynamics read
$d\bar{x}(t)=-\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N}.$
By consistency condition, the approximate average trading rate process it induces should satisfy
$\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}=\bar{Q}_{t},\quad t\in[0,T].$
Using the relationship between $\bar{P}(t),B(t)$ and $\bar{Q}(t)$, it is equivalent that the following equation system holds
$\left\{\begin{aligned} &d{\bar{P}(t)}=\alpha(\beta-\bar{P}(t)-\bar{Q}(t))dt,\quad\bar{P}(0)=p_{0},\\ &B^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{P}(t)+B(t))}{c}=0,\quad B(T)=-\gamma\zeta,\\ &d\bar{x}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N},\\ &\bar{Q}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}.\end{aligned}\right$.
By introducing matrix notation, we obtain an alternative formulation, which requires the following process
$\frac{d}{dt}\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix}=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\par\par\end{pmatrix}\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix}+\begin{pmatrix}\eta\kappa\\ 0\\ \alpha\beta\par\end{pmatrix}$, (2.13)
to hold true simultaneously with the following initial or terminal conditions
$\begin{pmatrix}B(T),\ \bar{x}(0),\ \bar{P}(0)\end{pmatrix}^{\top}=\begin{pmatrix}-\zeta\gamma,\ \bar{x}_{0}^{N},\ p_{0}\end{pmatrix}^{\top}$. (2.14)
The existence of solutions to (2.13) and (2.14) is crucial, for it determines whether or not the requirement in step 3 can be satisfied. The main difficulty in determining whether this equation has a solution lies in the fact that its boundary values are specified at different points in time. Therefore, when investigating the existence of its solution, we first analyze the properties of solutions given initial values at the same point, which will be presented in subsection 2.1 and subsection 2.1. They can be regarded as a preparation for the subsection 2.1, which proves that, under an assumption to be proposed (i.e. Assumption $\boldsymbol{(A_{1})}$), the existence and uniqueness of solutions to (2.13) and (2.14) hold true.
Proposition 2.3.
[…]
subsection 2.1 shows that the solution for (2.13) is uniquely determined by $B(0)=b_{0}$ once the conditions $\bar{x}(0)=\bar{x}_{0}^{N}$ and $\bar{P}(0)=p_{0}$ are given. Therefore, it remains to adjust the value of $b_{0}$ so that $B(T)=-\gamma\zeta$ holds true.
To this end, we construct the following process.
$\frac{d}{dt}\begin{pmatrix}B_{1}(t)\\ \bar{x}_{1}(t)\\ \bar{P}_{1}(t)\end{pmatrix}=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}B_{1}(t)\\ \bar{x}_{1}(t)\\ \bar{P}_{1}(t)\end{pmatrix},\forall t\in[0,T];\begin{pmatrix}B_{1}(0)\\ \bar{x}_{1}(0)\\ \bar{P}_{1}(0)\end{pmatrix}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$. (2.15)
Proposition 2.4.
There exists a unique solution to (2.15).
Now, we propose Assumption $\boldsymbol{(A_{1})}$, and we claim that under this assumption, (2.13), combined with (2.14), has a solution.
Assumption $\boldsymbol{(A_{1}):}$ The solution to (2.15) satisfies $B_{1}(T)\neq 0$.
Proposition 2.5.
If the unique solution to (2.15) satisfies $B_{1}(T)\neq 0$, (i.e. Assumption $\boldsymbol{(A_{1})}$ holds true), then (2.13), combined with (2.14), gives a unique solution.
We introduce another assumption, which places a constraint on the magnitude of the elements within the sets $\{\sigma_{i}\}_{1\leq i\leq N}$ and $\{x_{0}^{i}\}_{1\leq i\leq N}$.
Assumption $\boldsymbol{(A_{2}):}$ The sets $\{x_{0}^{i}\}_{i=1}^{N}$ and $\{\sigma_{i}\}_{i=1}^{N}$ are respectively defined on two fixed compact sets that are independent of $N$.
Now, we propose the main theorem for this section.
Theorem 2.6.
Assume that Assumption $\boldsymbol{(A_{1})}$ and $\boldsymbol{(A_{2})}$ hold true, and the unique solution to (2.13) and (2.14) is given by (B(t),x¯(t),P¯(t))(B(t),\ $\bar${x}(t),\ $\bar${P}($)^{\top}, t$hen the following system
$\left\{\begin{aligned} &d\hat{P}_{t}=\alpha(\beta-\hat{Q}_{t}-\hat{P}_{t})dt,\quad\hat{P}_{0}=p_{0},\\ &\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)\hat{X}_{t}^{i}+B(t)}{c},\\ &d\hat{X}_{t}^{i}=\hat{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad\hat{X}_{0}^{i}=x_{0}^{i},\\ &\hat{Q}_{t}=\frac{1}{N}\sum_{i=1}^{N}\hat{v}_{t}^{i}.\end{aligned}\right$. (2.16)
is an $\epsilon$-Nash equilibrium. Moreover, specific calculations show that $\epsilon_{N}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)}$ in (2.4).
Remark 2.
In its proof, the $O(1)$ numbers between adjacent lines may not be the same. This convention is also applied to the proofs in the cooperate games case. In addition, to present our derivation more clearly, we may not abbreviate specific important $O(1)$ numbers as $O(1)$, but instead denote them with specific letters.
In order to prove Theorem 2.6, we will introduce three lemmas. First, we introduce Lemma 2.1.
Lemma 2.1.
If $\hat{v}^{i}$, $\hat{X}^{i}$ and $\hat{Q}$ are given by (2.16), then there exists a positive constant $K^{\prime}$ such that
$\left\{\begin{aligned} &\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{X}_{t}^{i}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\int_{0}^{T}(\hat{v}_{t}^{i})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\hat{v}_{t}^{i})^{2}dt\right]\leq K^{\prime},\quad i=1,2,\ldots,N,\\ &\mathbb{E}\left[\sup_{0\leq t\leq T}|\overline{\hat{X_{t}}}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K^{\prime}\ ,\quad i=1,2,\ldots,N,\\ &\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\int_{0}^{T}\hat{Q}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{Q}_{t}|^{2}\right]\leq K^{\prime},\\ &|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq K^{\prime},\quad i=1,2,\ldots,N,\end{aligned}\right$.
where $\overline{\hat{X}}_{t}$ is the average of $\hat{X}_{t}^{i}$, from $i=1$ to $N$.
The proofs of all lemmas in this paper have been left to Appendix A.
To show that the system indeed induces an $\epsilon$-Nash equilibrium, we will show that for any agent $i$, its expected cost $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ does not decrease much if it changes its strategy unilaterally. Specifically, for agent $i(i=1,\ldots,N)$, we consider the case in which it adopts any admissible strategy $v^{i}\in\mathcal{A}$, while the strategies of all other agents remain fixed as $\hat{\boldsymbol{v}}^{-i}$. Then, agent $i$’s commodity evolves as follows
$dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i},$
and the evolution process for the price changes to
$d{P}_{t}=\alpha\bigg{(}\beta-\frac{v_{t}^{i}}{N}-\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}-{P}_{t}\bigg{)}dt,\quad{P}_{0}=p_{0}.$
To prove Theorem 2.6, we will compare the expected cost $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ with $J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$. Our aim is to prove (2.4) with $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$, or equivalently, for $i=1,\ldots,N$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)},\quad\forall\ v^{i}\in\mathcal{A}.$ (2.17)
In order to prove (2.17) , we need coercive condition, whose detailed statement is provided in Lemma 2.3. Prior to that, we present Lemma 2.2, which will facilitate a clear proof of Lemma 2.3.
Lemma 2.2.
For $N\in\mathbb{N}^{+}$, we construct the process
$\frac{d}{dt}\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix}=\begin{pmatrix}0&0\\ 0&{-\alpha}\end{pmatrix}\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix}+\begin{pmatrix}1\\ -\frac{\alpha}{N}\end{pmatrix}v_{t},\ \quad X_{0}^{*}=0,\ P_{0}^{*}=0.$
We claim that for any given $\epsilon_{1}^{*}>0$, $\epsilon_{2}^{*}>0$ and $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, we have
$\mathbb{E}\left[\int_{0}^{T}v_{t}(\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t})dt\right]\geq 0,\quad\forall v\in\mathcal{A}.$ (2.18)
Next, we introduce the coercive condition.
Lemma 2.3.
There exists a constant $C_{0}$, which is independent of $N$, such that the coercive condition
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq\mathbb{E}\left[\int_{0}^{T}\Big{(}\frac{\eta}{4}(X_{t}^{i}-\kappa)^{2}+\frac{c}{4}(v_{t}^{i})^{2}\Big{)}dt\right]+C_{0}$
is satisfied for $N>\frac{4(\sqrt{2}+1)}{c}$.
Now, we return to our proof of Theorem 2.6.
Proof of Theorem 2.6.
From Lemma 2.1, we have
$|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq K^{\prime}.$
For $\forall N>\frac{4(\sqrt{2}+1)}{c}$, if $\frac{c}{4}\|v^{i}\|^{2}+C_{0}\geq K^{\prime}$, then it follows from Lemma 2.3 that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}).$
This will make (2.17) hold. Otherwise, $\|v^{i}\|^{2}$ is bounded by $\frac{4(K^{\prime}-C_{0})}{c}$. Therefore, it suffices to prove (2.17) under the case where $\|v^{i}\|$ is bounded by an $O(1)$ constant.
As mentioned above, $\bar{P}(t)$ is constructed for approximating $P_{t}$ and $\hat{P}_{t}$. This inspires us to write the difference of the expected cost as
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$
$=\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right]-\mathbb{E}\left[\int_{0}^{T}(L(\hat{X}_{t}^{i},\hat{v}_{t}^{i})+\bar{P}(t)\hat{v}_{t}^{i})dt+\Psi(\hat{X}_{T}^{i})\right]$
$\quad+\mathbb{E}\Big{[}\int_{0}^{T}v_{t}^{i}(P_{t}-\bar{P}(t))dt\Big{]}-\mathbb{E}\Big{[}\int_{0}^{T}\hat{v}_{t}^{i}(\hat{P}_{t}-\bar{P}(t))dt\Big{]}$
$:=I_{1}+I_{2}-I_{3},$
where
$I_{1}$ $=\mathbb{E}\left[\int_{0}^{T}(L({X}_{t}^{i},{v}_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi({X}_{T}^{i})-\left(\int_{0}^{T}(L(\hat{X}_{t}^{i},\hat{v}_{t}^{i})+\bar{P}(t)\hat{v}_{t}^{i})dt+\Psi(\hat{X}_{T}^{i})\right)\right],$
$I_{2}$ $=\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}(P_{t}-\bar{P}(t))dt\right],$
$I_{3}$ $=\mathbb{E}\left[\int_{0}^{T}\hat{v}_{t}^{i}(\hat{P}_{t}-\bar{P}(t))dt\right].$
To finish the proof, it suffices to show that in the case where $\|v^{i}\|$ is bounded by an $O(1)$ number,
$I_{1}\geq 0,\quad I_{2}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)},\quad I_{3}=O\Big{(}\frac{1}{\sqrt{N}}\Big{)}.$
The first part is easy. We have proved that $\hat{v}_{t}^{i}$ is optimal for minimizing
$\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right].$
This gives $I_{1}\geq 0$.
The second and the third part is based on the fact that, when $\|v^{i}\|$ is bounded by an $O(1)$ number, $|I_{2}|$ and $|I_{3}|$ are bounded by $O(\frac{1}{\sqrt{N}})$ numbers. Specifically, we calculate both the difference between $\hat{P}_{t}$ and $\bar{P}(t)$ and that between $\bar{x}(t)$ and $\overline{\hat{X}}_{t}$. Direct calculation yields
$d\overline{\hat{X}}_{t}=-\frac{\bar{P}(t)+2a(t)\overline{{\hat{X}}}_{t}+B(t)}{c}dt+\frac{1}{N}\sum_{k=1}^{N}\sigma_{k}dW_{t}^{k},\quad\overline{\hat{X}}_{0}=\bar{x}_{0}^{N},$
$d\bar{x}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N},$
$\hat{P}_{t}-\bar{P}(t)=\frac{2\alpha}{c}\int_{0}^{t}e^{-\alpha(t-s)}a(s)\big{(}\overline{\hat{X_{s}}}-\bar{x}(s)\big{)}ds.$
It follows from Gronwall Lemma and Assumption $\boldsymbol{(A_{2})}$ that
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\bar{x}(t)-\overline{\hat{X}}_{t}\right|^{2}\right]\leq\mathbb{E}\left[O(1)\frac{1}{N^{2}}\left(\sqrt{\sum_{k=1}^{N}\sigma_{k}^{2}}\right)^{2}\right]=O\left(\frac{1}{N}\right),$
which implies (from the Cauchy-Schwartz inequality)
$\|\hat{P}-\bar{P}\|^{2}$ $\leq O(1)\mathbb{E}\left[\int_{0}^{T}\bigg{(}\int_{0}^{t}\left(e^{-\alpha(t-s)}\right)^{2}ds\bigg{)}\bigg{(}\int_{0}^{t}\left(\overline{\hat{X}}_{s}-\bar{x}(s)\right)^{2}ds\bigg{)}dt\right]$
$\leq O(1)\left(\int_{0}^{T}\bigg{(}\int_{0}^{t}\left(e^{-\alpha(t-s)}\right)^{2}ds\bigg{)}dt\right)\mathbb{E}\left[\int_{0}^{T}\left(\overline{\hat{X}}_{s}-\bar{x}(s)\right)^{2}ds\right]$
$=O\left(\frac{1}{N}\right).$
Then, it is convenient and useful for us to compare $P_{t}$ with $\hat{P}_{t}$,
$P_{t}-\hat{P}_{t}=-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}(v_{s}^{i}-\hat{v}_{s}^{i})ds.$
We deduce from the Cauchy-Schwartz inequality ,
$\|P-\hat{P}\|^{2}\leq\frac{1}{N^{2}}\left(\int_{0}^{T}dt\int_{0}^{t}(e^{-\alpha(t-s)}\alpha)^{2}ds\right)\mathbb{E}\left[\int_{0}^{T}(v_{s}^{i}-\hat{v}_{s}^{i})^{2}ds\right]=O\bigg{(}\frac{1}{N^{2}}\bigg{)}.$
The last equality is based on the fact that $\|\hat{v}^{i}\|$ and $\|v^{i}\|$ are all bounded by $O(1)$ numbers. The above estimates imply that
$\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),\quad\|P-\hat{P}\|=O\bigg{(}\frac{1}{N}\bigg{)},\quad\|P-\bar{P}\|\leq\|P-\hat{P}\|+\|\hat{P}-\bar{P}\|\leq O\left(\frac{1}{\sqrt{N}}\right),$
where the last result follows from the triangle inequality. By using the Cauchy-Schwartz inequality again
$|I_{2}|\leq\|v^{i}\|\cdot\|P-\bar{P}\|=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)},$
$|I_{3}|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}-\bar{P}\|=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
We complete the proof for the second and the third part. Finally, we conclude that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
Hence, the proof is complete. ∎
2.2 The Cooperate Games
In this subsection, we are looking for a set of strategies that achieve an $\epsilon$-optimal expected social cost. Since our objective is to minimize $J_{soc}$, it is necessary for us to comprehensively measure the impact of agent $i$’s strategy on $J_{soc}$. To this end, We decompose the price process $P$ into components that are influenced and not influenced by agent $i$, respectively, and proceed to analyze the effect of $v^{i}$ on $J_{soc}$.
Our approach can be outlined in the following steps.
  1. 1.
    Show that under the optimal strategies of $J_{soc}$, agent $i$’s strategy must be the minimizer of the following auxiliary control problem:
    $\check{J}_{i}(v^{i})$ $:=\mathbb{E}\left[\int_{0}^{T}\bigg{(}\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}P_{t}-\frac{u_{t}^{i}}{N}\sum_{k=1,k\neq i}^{N}\check{v}_{t}^{k}\bigg{)}dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right],$ (2.19)
    subject to
    $\left\{\begin{aligned} &\frac{dP_{t}}{dt}=\alpha\bigg{(}\beta-\frac{1}{N}\sum_{k\neq i}^{N}\check{v}_{t}^{k}-\frac{1}{N}v_{t}^{i}-P_{t}\bigg{)},\quad P_{0}=p_{0},\\ &dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i},\\ &u_{t}^{i}=\int_{0}^{t}e^{-\alpha(t-s)}\alpha v_{s}^{i}ds.\end{aligned}\right$. (2.20)
  2. 2.
    We approximate $\frac{1}{N}\sum_{k\neq i}\check{v}_{t}^{k}$ with a continuous function $(\bar{q}(t))_{0\leq t\leq T}$. Then, we derive a continuous function $(\bar{p}(t))_{0\leq t\leq T}$, which can be viewed as the approximation of price process $(P_{t})_{t\in[0,T]}$. It evolves according to
    $d\bar{p}(t)=\alpha(\beta-\bar{p}(t)-\bar{q}(t))dt,\quad\bar{p}(0)=p_{0}.$ (2.21)
  3. 3.
    For $i=1,\ldots,N$, by applying the approximation introduced above, we obtain the following auxiliary objective function for agent $i$ from $\check{J}_{i}$,
    $J^{\prime}_{i}(v^{i};\bar{p},\bar{q})=\mathbb{E}\left[\int_{0}^{T}\Big{(}\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)-\bar{q}(t)u_{t}^{i}\Big{)}dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right],$ (2.22)
    subject to $(X^{i},u^{i})$ with (2.20).
  4. 4.
    Solve control problem $J^{\prime}_{i}$, obtain the optimal control strategies, and construct the consistence condition which is satisfied by $\bar{q}$. Finally, we show that the strategies derived from the solution to this consistence condition indeed has asymptotic social optimality.
Our first step is to evaluate precisely how the strategy of a certain agent affects the expected social cost $J_{soc}$. To do this, we divide $J_{soc}$ into three parts, one of them being closely related to its own strategy, the rest being independent of its control. To be more specific, we state that:
Theorem 2.7.
If the strategies $\{\check{v}_{t}^{i}\}_{i=1}^{N}$ optimize $J_{soc}$, then for any $i=1,\ldots,N$, $\check{v}_{t}^{i}$ is the solution for minimizing control problem (2.19)-(2.20) when other agents’ strategies are given by $\check{v}^{-i}$.
Proof.
We divide $J_{soc}$ into different parts. The objective is to evaluate the overall impact of agent $i$’s strategy on $J_{soc}$. Let $J_{soc}=Y_{1}^{i}+Y_{2}^{i}$, for $i=1,\ldots,N$, where
$Y_{1}^{i}:=\frac{1}{N}\sum_{k\neq i}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{k}-\kappa)^{2}+\frac{c}{2}(v_{t}^{k})^{2}\right)dt+\frac{\gamma}{2}\left(X_{T}^{k}-\zeta\right)^{2}\right],$
$Y_{2}^{i}:=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+\Big{(}\sum_{k=1}^{N}v_{t}^{k}\Big{)}P_{t}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right].$
It should be noted that the evolution of $P$ is only partially influenced by strategy $v^{i}$. Therefore, we solve from (2.2)
$P_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds,$
and apply $Q_{t}=\frac{1}{N}v_{t}^{i}+\frac{1}{N}\sum_{k\neq i}v_{t}^{k}$ to obtain $Y_{2}^{i}=Y_{3}^{i}+Y_{4}^{i}$, where
$Y_{3}^{i}$ $=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{k\neq i}v_{t}^{k}\right)\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\left(\sum_{k\neq i}v_{s}^{k}\right)ds\right)dt\right],$
$Y_{4}^{i}$ $=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}{(v_{t}^{i})}^{2}+v_{t}^{i}P_{t}\right)dt+\frac{\gamma}{2}\left(X_{T}^{i}-\zeta\right)^{2}\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{k\neq i}v_{t}^{k}\right)\left(-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds\right)dt\right].$
Observe that $J_{soc}=Y_{1}^{i}+Y_{3}^{i}+Y_{4}^{i}$, and $Y_{1}^{i}$, $Y_{3}^{i}$ are both independent of $v^{i}$. Combining with $Y_{4}^{i}=\frac{1}{N}\check{J}_{i}(v^{i})$, we conclude that if $(\check{v}_{t}^{1},...,\check{v}_{t}^{N})$ minimizes $J_{soc}$, then $\check{v}_{t}^{i}$ minimizes $\check{J}_{i}(v^{i})$ for $i=1,\ldots,N$. The proof is complete. ∎
Subsequently, we solve the optimal control problem (2.22). It follows from (2.20) that
$du_{t}^{i}=(-\alpha u_{t}^{i}+\alpha v_{t}^{i})dt,\quad u_{0}^{i}=0.$
By dynamic programming principle, the value function
$V_{i}(t,x,u)=\inf_{v^{i}\in\mathcal{A}}\mathbb{E}\left[\int_{t}^{T}\Big{(}\frac{\eta}{2}(X_{s}^{i}-\kappa)^{2}+\frac{c}{2}(v_{s}^{i})^{2}-\bar{q}(s)u_{s}^{i}+v_{s}^{i}\bar{p}(s)\Big{)}ds+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
satisfies the following HJB equation
$\frac{\partial V_{i}}{\partial t}+\inf_{v^{i}\in\mathcal{A}}\left\{\frac{\partial V_{i}}{\partial x}v_{t}^{i}+\frac{\partial V_{i}}{\partial u}(\alpha v_{t}^{i}-\alpha u)+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)\right\}-\bar{q}(t)u+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}V_{i}}{\partial^{2}x}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
and the terminal condition $V_{i}(T,x,u)=\frac{\gamma}{2}(x-\zeta)^{2}$. We consider the ansatz
$V_{i}(t,x,u)=a(t)x^{2}+b(t)x+l(t)u+f_{i}(t).$
The HJB equation then reduces to
$a^{\prime}(t)x^{2}+b^{\prime}(t)x+f_{i}^{\prime}(t)+l^{\prime}(t)u+\inf_{v^{i}\in\mathcal{A}}\Big{\{}(2a(t)x+b(t))v_{t}^{i}+\alpha l(t)v_{t}^{i}-\alpha ul(t)$
$\quad+\frac{c}{2}{(v_{t}^{i})}^{2}+v_{t}^{i}\bar{p}(t)\Big{\}}-\bar{q}(t)u+\sigma_{i}^{2}a(t)+\frac{\eta}{2}(x-\kappa)^{2}=0.$
It yields the exact form of the optimal control
$\check{v}_{t}^{i}=-\frac{2a(t)\check{X}_{t}^{i}+b(t)+\bar{p}(t)+\alpha l(t)}{c}.$ (2.23)
Applying (2.23), we simplify the HJB equation into the following ODE system
$\qquad\qquad\qquad a^{\prime}(t)+\frac{\eta}{2}-\frac{2a^{2}(t)}{c}=0,$ $a(T)=\frac{\gamma}{2},$ (2.24)
$\qquad\qquad\qquad b^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{p}(t)+b(t)+\alpha l(t))}{c}=0,$ $b(T)=-\gamma\zeta,$ (2.25)
$\qquad\qquad\qquad f^{\prime}_{i}(t)+\sigma_{i}^{2}a(t)+\frac{\eta\kappa^{2}}{2}-\frac{(b(t)+\bar{p}(t)+\alpha l(t))^{2}}{2c}=0,$ $f_{i}(T)=\frac{\gamma\zeta^{2}}{2},$ (2.26)
$\qquad\qquad\qquad l^{\prime}(t)-\alpha l(t)-\bar{q}(t)=0,$ $l(T)=0.$ (2.27)
We state that (2.24),(2.25),(2.26) and (2.27) give a unique solution established on $[0,T]$. To prove this, we may argue as follows. First, (2.24) gives a unique solution on $[0,T]$, as a result of subsection 2.1. Then, similar as in the proof of subsection 2.1, (2.27) and (2.25) both give a unique solution on $[0,T]$. Finally, the existence and uniqueness of solutions to (2.26) is obvious, once we integrate both sides of (2.26) from $t$ to $T$. We will not verify the the optimality of the solution given by (2.23), for the optimality property is not necessary for us to arrive at the results in Theorem 2.11.
As in the discussion of the non-cooperate games section, if the agents adopt the strategies introduced in (2.23), the average of $X_{t}^{i}$, abbreviated as $\overline{X}_{t}$, can be approximated by a deterministic process $\bar{x}$, whose dynamics satisfy
$d\bar{x}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N}.$
This implies that, the requirement in step 3 is equivalent to $\bar{q}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}$. By integrating the process (2.21),(2.25), (2.27) and the above evolution law for $\bar{q}$, we observe that this is also equivalent to ensure that the following dynamics
$\frac{d}{dt}\begin{pmatrix}\bar{p}(t)\\ \bar{x}(t)\\ b(t)\\ l(t)\end{pmatrix}=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}(t)\\ \bar{x}(t)\\ b(t)\\ l(t)\end{pmatrix}+\begin{pmatrix}\alpha\beta\\ 0\\ \eta\kappa\\ 0\end{pmatrix}$, (2.28)
together with the corresponding initial or terminal conditions,
$\begin{pmatrix}\bar{p}(0),\ \bar{x}(0),\ b(T),\ l(T)\end{pmatrix}^{\top}=\begin{pmatrix}p_{0},\ \overline{x}_{0}^{N},\ -\gamma\zeta,\ 0\end{pmatrix}^{\top}$, (2.29)
holds true simultaneously.
Again, it is crucial that (2.28), together with (2.29), indeed gives a solution on $[0,T]$. Similar as in the previous case, the main difficulty in determining the existence of this solution is that the boundary values are specified at different points. By this observation, we first construct two processes whose initial values are only specified at $t=0$ and study their properties.
Proposition 2.8.
Equation systems
$\frac{d}{dt}\begin{pmatrix}\bar{p}_{1}(t)\\ \bar{x}_{1}(t)\\ {b}_{1}(t)\\ {l}_{1}(t)\end{pmatrix}$ $=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}_{1}(t)\\ \bar{x}_{1}(t)\\ {b}_{1}(t)\\ {l}_{1}(t)\end{pmatrix}$, (2.30)
and
$\frac{d}{dt}\begin{pmatrix}\bar{p}_{2}(t)\\ \bar{x}_{2}(t)\\ {b}_{2}(t)\\ {l}_{2}(t)\end{pmatrix}$ $=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}_{2}(t)\\ \bar{x}_{2}(t)\\ {b}_{2}(t)\\ {l}_{2}(t)\end{pmatrix}$, (2.31)
with $(\bar{p}_{1}(0),\ \bar{x}_{1}(0),\ {b}_{1}(0),\ {l}_{1}(0))^{\top}=(0,0,1,0)^{\top}$ and $(\bar{p}_{2}(0),\ \bar{x}_{2}(0),\ {b}_{2}(0),\ {l}_{2}(0))^{\top}=(0,0,0,1)^{\top}$, give a unique solution, denoted by $(\phi_{1}^{*}(t))_{t\in[0,T]}=((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $(\phi_{2}^{*}(t))_{t\in[0,T]}$ $((\bar{p}_{2}(t),\ \bar{x}_{2}(t),\ b_{2}(t),\ l_{2}(t))^{\top})_{t\in[0,T]}$ respectively.
Similar as in the non-cooperate case, we will show that, the solution to (2.28), if given initial values specified only at $t=0$, must give a unique solution.
Proposition 2.9.
For $\forall b_{0},\ l_{0}\in\mathbb{R}$, the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a unique solution that satisfies (2.28), denoted by
$(\phi_{b_{0},l_{0}}^{*}(t))_{t\in[0,T]}=((\bar{p}_{b_{0},l_{0}}(t),\ \bar{x}_{b_{0},l_{0}}(t),\ b_{b_{0},l_{0}}(t),\ l_{b_{0},l_{0}}(t))^{\top})_{t\in[0,T]}.$
subsection 2.2 states that, given $\bar{p}(0)=p_{0}$ and $\bar{x}(0)=\bar{x}_{0}^{N}$, the solution to (2.28) is uniquely determined by $b(0)$ and $l(0)$. Now, we propose a sufficient condition for the existence and uniqueness for (2.28) and (2.29), as shown in Assumption $\boldsymbol{(A_{3})}$.
Assumption $\boldsymbol{(A_{3}):}$ The processes $((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $((\bar{p}_{2}(t),\ \bar{x}_{2}(t)$, $b_{2}(t),$ $l_{2}(t))^{\top})_{t\in[0,T]}$, which are given by (2.30) and (2.31) respectively, satisfy that $({b}_{1}(T),\ {l}_{1}(T))^{\top}$, $({b}_{2}(T),\ {l}_{2}(T))^{\top}$ are linearly independent vectors.
Proposition 2.10.
If Assumption $\boldsymbol{(A_{3})}$ holds true, then the existence and uniqueness of (2.28) and (2.29) hold true.
As in other MFG works, We need coercive conditions to prevent the absence of an optimal or $\epsilon$-optimal control. However, this condition may not hold for all positive numbers $c$ and $\eta$. This drives us to propose the following assumption to ensure that the coercive condition holds true.
Assumption $\boldsymbol{(A_{4}):}$ The parameters $c$ and $\eta$ satisfy that $(c-2)\eta>\alpha^{2}$.
Remark 3.
To see why the assumption is necessary, consider the following case. If $v_{t}^{i}$ remains being a large constant, then $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}(v_{t}^{i}P_{t})dt\right]$ grows into a very small negative number that may exceed the influence of the positive terms $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\left(\frac{c}{2}{(v_{t}^{i})}^{2}+\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}\right)dt\right]$ when $c$ and $\eta$ are sufficiently small. Therefore, $J_{soc}$ can be arbitrarily small as the norm of the control increases, which makes the coercive condition fail, and the discussion of the problem meaningless. Note that the coefficient of the term $v_{t}^{i}P_{t}$ is 1 within $J_{soc}$, we posit that $\frac{c}{2}>1$ and proper lower bound for $\eta$ would suffice to guarantee coercivity; this will be rigorously shown in Lemma 2.5 below.
Given the aforementioned assumptions, we propose the process as shown in Theorem 2.11, and demonstrate that it provides a set of $\epsilon$-optimal strategies.
Theorem 2.11.
Assume that Assumption $\boldsymbol{(A_{2})}$, $\boldsymbol{(A_{3})}$ and $\boldsymbol{(A_{4})}$ hold, and the unique solution for (2.28) and (2.29) is $(\bar{p}(t),\ \bar{x}(t),\ b(t),\ l(t))^{\top}$. We construct the process
$\left\{\begin{aligned} &d\check{X}_{t}^{i}=\check{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad\check{X}_{0}^{i}=x_{0}^{i},\\ &\check{v}_{t}^{i}=-\frac{2a(t)\check{X}_{t}^{i}+{b}(t)+\bar{p}(t)+\alpha{l}(t)}{c},\\ &\check{Q}_{t}=\frac{1}{N}\sum_{i=1}^{N}\check{v}_{t}^{i},\\ &d\check{P}_{t}=\alpha(\beta-\check{P}_{t}-\check{Q}_{t})dt,\quad\check{P}_{0}=p_{0}.\end{aligned}\right$. (2.32)
Then, we claim that the set of strategies $\{\check{v}^{i}\}_{1\leq i\leq N}$ is a set of $\epsilon$-optimal strategies for $J_{soc}$. Moreover, specific calculation shows $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$ in (2.5).
Assume there exists an alternative set of admissible control strategies $\boldsymbol{v}=\{v^{i}\}_{1\leq i\leq N}$, which differs from $\check{\boldsymbol{v}}=\{\check{v}^{i}\}_{1\leq i\leq N}$. The corresponding processes it induces evolve as follows
$\left\{\begin{aligned} &d{X_{t}^{i}}={v_{t}^{i}}dt+\sigma_{i}dW_{t}^{i},\quad X_{t}^{i}=x_{0}^{i},\\ &Q_{t}=\frac{1}{N}\sum_{i=1}^{N}{v_{t}^{i}},\\ &dP_{t}=\alpha(\beta-P_{t}-Q_{t})dt,\quad P_{0}=p_{0}.\end{aligned}\right$.
Our objective is to show that (2.5) holds true, or equivalently
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}})\geq-\epsilon_{N},$ (2.33)
where the $\epsilon_{N}$ term is an $O(\frac{1}{\sqrt{N}})$ number. To do this, we introduce two lemmas. Lemma 2.5 is proposed to characterize the coercive condition. However, to facilitate a clear demonstration of the coercivity, we begin by presenting Lemma 2.4.
Lemma 2.4.
We construct the process
$x_{t}=\begin{pmatrix}\tilde{X}_{t}\\ \tilde{P}_{t}\end{pmatrix},\quad\frac{d}{dt}x_{t}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}x_{t}+\begin{pmatrix}1\\ -\alpha\end{pmatrix}v_{t},\quad x_{0}=\begin{pmatrix}0\\ 0\end{pmatrix}$.
Now, we claim that, for any stochastic control $v\in\mathcal{A}$, the following result
$\mathbb{E}\left[\int_{0}^{T}v_{t}\Big{(}v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}\Big{)}dt\right]\geq 0$
holds true.
We now proceed to the rigorous formulation of the coercivity condition, as shown in Lemma 2.5.
Lemma 2.5.
Assume that Assumptions $\boldsymbol{(A_{2})}$ and $\boldsymbol{(A_{4})}$ hold true. Then there exists constants $\epsilon>0,\ \epsilon^{\prime}>0$, and $C\in\mathbb{R}$, all independent of $N$, such that
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}({\epsilon}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2})dt\right]+C.$
Up to this point, we have done the preparations, so we proceed to the proof of Theorem 2.11.
Proof of Theorem 2.11.
Denote the difference in each process by
$\Delta X_{t}^{i}=X_{t}^{i}-\check{X}_{t}^{i},\quad\Delta v_{t}^{i}=v_{t}^{i}-\check{v}_{t}^{i},\quad\Delta P_{t}=P_{t}-\check{P}_{t},\quad\Delta Q_{t}=Q_{t}-\check{Q}_{t},\quad\Delta u_{t}^{i}=u_{t}^{i}-\check{u}_{t}^{i},$
where
$du_{t}^{i}=-\alpha u_{t}^{i}dt+\alpha v_{t}^{i}dt,\quad u_{0}^{i}=0,$
$d\check{u}_{t}^{i}=-\alpha\check{u}_{t}^{i}dt+\alpha\check{v}_{t}^{i}dt,\quad\check{u}_{0}^{i}=0.$
and $(\check{X}_{t}^{i},\ \check{P}_{t},\ \check{Q}_{t})$ are given by (2.32). Within these notions, the variation in $J_{soc}$ can be written as
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}})$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(\check{X}_{t}^{i}+\Delta X_{t}^{i}-\kappa\right)^{2}-\frac{\eta}{2}\left(\check{X}_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}\left(\check{v}_{t}^{i}+\Delta v_{t}^{i}\right)^{2}-\frac{c}{2}({\check{v}_{t}^{i})}^{2}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\check{v}_{t}^{i}+\Delta{v_{t}^{i}})(\check{P}_{t}+\Delta{P}_{t})-\check{v}_{t}^{i}\check{P}_{t}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\frac{\gamma}{2}\left(\check{X}_{T}^{i}+\Delta X_{T}^{i}-\zeta\right)^{2}-\frac{\gamma}{2}\left(\check{X}_{T}^{i}-\zeta\right)^{2}\right].$ (2.34)
We transform the difference in terminal value into its integration form by using Ito’s formula:
$\mathbb{E}\bigg{[}\frac{\gamma}{2}\left(\check{X}_{T}^{i}+\Delta X_{T}^{i}-\zeta\right)^{2}-\frac{\gamma}{2}\left(\check{X}_{T}^{i}-\zeta\right)^{2}\bigg{]}$
$=\mathbb{E}\left[\int_{0}^{T}\left(a^{\prime}(t)(\Delta X_{t}^{i})^{2}+2a^{\prime}(t)\Delta X_{t}^{i}\check{X}_{t}^{i}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(2a(t)\Delta X_{t}^{i}\check{v}_{t}^{i}+2a(t)\Delta v_{t}^{i}\check{X}_{t}^{i}+{b}^{\prime}(t)\Delta X_{t}^{i}+{b}(t)\Delta v_{t}^{i}\right)dt\right].$ (2.35)
Plugging (2.35) into (2.34), we write
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}}):=I_{1}+I_{2},$
where
$I_{1}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\bigg{[}\int_{0}^{T}\left((\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)+\Delta{X_{t}^{i}}\check{X}_{t}^{i}\left(\eta+2a^{\prime}(t)-\frac{(2a(t))^{2}}{c}\right)\right)dt\bigg{]}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\Delta{X_{t}^{i}}\left(-\eta\kappa+{b}^{\prime}(t)-\frac{2a(t)(\bar{p}(t)+\alpha{l}(t)+{b}(t))}{c}\right)+\Delta v_{t}^{i}\left(-\alpha l(t)\right)\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}+\bar{q}(t)\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right]$
$=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)-\Delta v_{t}^{i}\alpha{l}(t)+\frac{c}{2}(\Delta v_{t}^{i})^{2}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}+\Delta v_{t}^{i}\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right],$ (2.36)
and
$I_{2}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\check{v}_{t}^{i}-\bar{q}(t))\Delta P_{t}+\Delta Q_{t}(\check{P}_{t}-\bar{p}(t)\right)dt\right].$
We now provide a more detailed explanation of how we derive $I_{1}$ and what $I_{2}$ characterizes. To obtain $I_{1}$, we first approximate $\sum_{i=1}^{N}\check{v}_{t}^{i}\Delta P_{t}$ with $\sum_{i=1}^{N}\overline{q}(t)\Delta P_{t}$ and approximate $\sum_{i=1}^{N}\check{P}_{t}\Delta v_{t}^{i}$ with $\sum_{i=1}^{N}\overline{p}(t)\Delta v_{t}^{i}$ when calculating the integral of $\sum_{i=1}^{N}(\Delta P_{t}\check{v}_{t}^{i}+\check{P}_{t}\Delta v_{t}^{i})$. Then, we apply (2.35) to transform the terminal terms into the integral form. Following this, we use (2.32) to eliminate the $\check{v}_{t}^{i}$ term (i.e. turn it into the expression of $\check{X}_{t}^{i}$ and the continuous functions). Finally, we simply integrate the same terms to obtain $I_{1}$. Obviously, it characterizes how much approximately $J_{soc}$ increases after a deviation of the proposed process is observed. The second part, $I_{2}$, explicitly measures the error caused by this approximation.
Working exactly as in the proof of Lemma 2.1, we obtain that there exists an $O(1)$ constant, which we set as $K^{\prime\prime}$, that satisfies
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\check{X}_{t}^{i}|^{2}\right]\vee\mathbb{E}\left[\sup_{0\leq t\leq T}|\check{Q}_{t}|^{2}\right]\leq K^{\prime\prime},\ \mathbb{E}\left[\int_{0}^{T}({\check{v}_{t}^{i}})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\check{v}_{t}^{i})^{2}dt\right]\leq K^{\prime\prime},\ i=1,\ldots,N,$
and
$|J_{i}(\check{v}^{i},\check{\boldsymbol{v}}^{-i})|\leq K^{\prime\prime},\quad i=1,2,\ldots,N.$
This leads to
$J_{soc}(\check{\boldsymbol{v}})\leq K^{\prime\prime}.$
From the coercive condition given by Lemma 2.5, we conclude that there exists an $O(1)$ constant $M_{1}>0$ such that, if $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]>M_{1}$ is satisfied, then we have
$J_{soc}(\boldsymbol{v})\geq J_{soc}(\check{\boldsymbol{v}}).$
This inequality would give us the desired result. Therefore, it suffices to prove (2.33) under the case where $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]\leq M_{1}$. Indeed, under this case, we will demonstrate that
$I_{1}\geq 0,\quad I_{2}=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
First, we show that the first part holds. The next step relies on the following identity
$\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=0.$ (2.37)
We will then demonstrate how it is proved. Ito’s formula, together with l{l}’s evolution process displayed in (2.27), implies that
$0$ $=\mathbb{E}\left[{l}(T)\Delta u_{T}^{i}\right]=\mathbb{E}\left[\int_{0}^{T}\left({l}^{\prime}(t)\Delta u_{t}^{i}+{l}(t)(-\alpha\Delta u_{t}^{i}+\alpha\Delta v_{t}^{i})\right)dt\right]$
$=\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta u_{t}^{i}dt+\alpha{l}(t)\Delta v_{t}^{i}\right)dt\right].$ (2.38)
$\Delta{P_{t}}$ and $\Delta u_{t}^{i}$ evolve as follows,
$d\Delta{P_{t}}=\alpha(-\Delta{P_{t}}-\Delta Q_{t})dt,\quad\Delta{P_{0}}=0,$ (2.39)
$d\Delta u_{t}^{i}=\alpha(-\Delta u_{t}^{i}+\Delta v_{t}^{i})dt,\quad\Delta u_{0}^{i}=0.$ (2.40)
Note that $\Delta Q_{t}$ is the average of $\Delta v_{t}^{i}$, it is easily checked from (2.39) and (2.40) that
$\frac{1}{N}\sum_{i=1}^{N}\Delta u_{t}^{i}=-\Delta P_{t}.$ (2.41)
This gives
$\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(-\bar{q}(t)\Delta u_{t}^{i}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=0,$
where the first and the second equation are derived from (2.41) and (2.2) respectively. This proves (2.37). By a simple calculation, (2.37) simplifies (2.36) into
$I_{1}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{i=1}^{N}(\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)+\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right].$
Using Ito’s formula again, we obtain
$\mathbb{E}\left[a(T)(\Delta X_{T}^{i})^{2}\right]=\mathbb{E}\left[\int_{0}^{T}\left(a^{\prime}(t)(\Delta X_{t}^{i})^{2}+2a(t)\Delta X_{t}^{i}\Delta v_{t}^{i}\right)dt\right].$
This further simplifies the expression of $I_{1}$ into
$I_{1}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(\Delta{X_{t}^{i}})^{2}+\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}\right)dt+\frac{\gamma}{2}(\Delta X_{T}^{i})^{2}\right].$
We introduce the following notation
$\overline{\Delta X_{t}}:=\frac{1}{N}\sum_{i=1}^{N}\Delta X_{t}^{i}.$
Note that
$\left\{\begin{aligned} &d\overline{\Delta X_{t}}=\Delta Q_{t}dt,\\ &d\Delta P_{t}=-\alpha\Delta P_{t}dt-\alpha\Delta Q_{t}dt,\\ &\overline{\Delta X_{0}}=0,\quad\Delta P_{0}=0.\end{aligned}\right$.
$(\overline{\Delta X_{t}},\ \Delta P_{t})^{\top}$ follows the evolution process introduced in Lemma 2.4, so we conclude
$\mathbb{E}\left[\int_{0}^{T}\left(\Delta Q_{t}+\Delta P_{t}+\alpha\overline{\Delta X_{t}}\right){\Delta Q_{t}}dt\right]\geq 0.$
In order to use this inequality, we make use of Assumption $\boldsymbol{(A_{4})}$. Fundamental inequality implies
$\frac{\eta}{2}\left(\overline{\Delta X_{t}}\right)^{2}-\alpha\overline{\Delta X_{t}}\Delta Q_{t}+\frac{c-2}{2}\Delta Q_{t}^{2}\geq 0.$
As in the proof of Theorem 2.6, the convexity of $\frac{\eta}{2}x^{2}$ leads to
$I_{1}$ $\geq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(\overline{\Delta{X_{t}}}\right)^{2}+\frac{c}{2}\left(\Delta Q_{t}\right)^{2}+\Delta v_{t}^{i}\Delta P_{t}\right)dt\right]$
$=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\Delta Q_{t}(\Delta P_{t}+\alpha\overline{\Delta X_{t}}+\Delta Q_{t})\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(\overline{\Delta X_{t}})^{2}-\alpha\overline{\Delta X_{t}}\Delta Q_{t}+\frac{c-2}{2}(\Delta Q_{t})^{2}\right)dt\right],$ (2.42)
where we apply Jensen’s inequality. Specifically, we substitute $\sum_{i=1}^{N}(\Delta X_{t}^{i})^{2}$ with the smaller (or equal) number $N\overline{\Delta X_{t}}^{2}$, and $\sum_{i=1}^{N}(\Delta v_{t}^{i})^{2}$ with $N\Delta Q_{t}^{2}$. Note that (2.42) gives a non-negative value, so
$I_{1}\geq 0.$
The proof for the first part is completed, and it is straightforward to prove the second part. Now, similar as in Theorem 2.6’s proof, we have
$\mathbb{E}\left[\int_{0}^{T}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\big{(}\check{v}_{t}^{i}-\bar{q}(t)\big{)}\bigg{)}^{2}dt\right]=O\Big{(}\frac{1}{N}\Big{)}.$
From the Cauchy-Schwartz inequality, the triangle inequality, and $\|Q\|^{2}\leq M_{1}$, we have
$\|\Delta Q\|^{2}\leq O(1)(\|Q\|+\|\check{Q}\|)^{2}\leq O(1).$
We solve from (2.39),
$\Delta P_{t}=-\alpha\int_{0}^{t}e^{-\alpha(t-s)}\Delta Q_{s}ds.$
Cauchy-Schwartz theorem implies that
$|\Delta P_{t}|^{2}\leq\alpha^{2}\|\Delta Q\|^{2}\left(\int_{0}^{t}e^{-2\alpha(t-s)}ds\right)\leq O(1)\|\Delta Q\|^{2},\quad\|\Delta P\|^{2}\leq O(1)\|\Delta Q\|^{2}.$
Thus,
$\frac{1}{N}\left|\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\big{(}\check{v}_{t}^{i}-\bar{q}(t)\big{)}\Delta P_{t}dt\right]\right|\leq\frac{1}{N}\|\Delta P\|\cdot\left\|\sum_{i=1}^{N}(\check{v}^{i}-\bar{q})\right\|\leq O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$ (2.43)
Proceeding in the same way as in Theorem 2.6, we get
$\|\check{P}-\bar{p}\|=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
Therefore,
$\bigg{|}\mathbb{E}\left[\int_{0}^{T}\Delta Q_{t}(\check{P}_{t}-\bar{p}(t))dt\right]\bigg{|}\leq\|\Delta Q\|\cdot\|\check{P}-\bar{p}\|\leq O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$ (2.44)
We deduce from (2.43) and (2.44) that
$I_{2}=O\bigg{(}\frac{1}{\sqrt{N}}\bigg{)}.$
The proof is complete.
3 Numerical experiments
[…]
Then, we compare the discrepancy between the approximation functions and the true values under the non-cooperate case and the cooperate case through numerical experiments. As shown below, 1(a) and 1(b) compare the evolution of $\bar{p}$ and $P$, $\bar{q}$ and $Q$ in the non-cooperative game setting, respectively. 2(a) and 2(b) present the same comparisons under the cooperative game setting. From the plot, we can see that in the non-cooperate case, both curves in 1(a) exhibit a decreasing trend, and both curves in 1(b) display a $U$-shaped trajectory. However, the cooperative strategies not only lead to a steeper initial decline in the price, but also cause $Q$ to decrease monotonically. Besides, the cooperative case yields more stable and smoother trajectories, especially for the $Q$-related dynamics. This suggests that cooperation among agents can lead to more efficient and consistent outcomes in price and the average trading rate. Although the shapes of the $P$ and $Q$ curves differ between the figures in both cases, these figures all demonstrate that the mean-field approximation performs very well in both cooperative and non-cooperative settings. This property could be valuable in practical applications like smart grids or large-scale decentralized systems.
(a) Curves of $\bar{P}$ and $P$ in the non-cooperate case
(b) Curves of $\bar{Q}$ and $Q$ in the non-cooperate case
Figure 1: Non-cooperate case: evolution of $\bar{P},P$ and $\bar{Q},Q$
(a) Curves of $\bar{p}$ and $P$ in the cooperate case
(b) Curves of $\bar{q}$ and $Q$ in the cooperate case
Figure 2: Cooperate case: evolution of $\bar{p},P$ and $\bar{q},Q$
As shown in Figure 3, we also compare the evolution of the control variable $v$ for the first 10 agents under both cases. In 3(a), which corresponds to the non-cooperative case, the trajectories of $v^{i}$ appear more tightly clustered, with relatively smaller deviations from one another over time. In contrast, 3(b), which shows the cooperative case, exhibits greater divergence in the trajectories of $v^{i}$. Overall, the figures illustrate how cooperation introduces more variability in individual control paths, while non-cooperation leads to more uniform behavior across agents. This reflects the fact that, in a cooperative setting, agents can adjust their behaviors more flexibly to improve overall system performance.
(a) Evolution of $v^{i}$ for agent 1 to 10 (the non-cooperate case)
(b) Evolution of $v^{i}$ for agent 1 to 10 (the cooperate case)
Figure 3: Comparison of the evolution process of $v^{i}$ under both cases
4 Conclusion
This paper has developed a strategic framework for smart grids with many agents using mean field game methodology. By analyzing both cooperative and non-cooperative scenarios under sticky prices and finite time horizons, we provide approaches to minimize expected losses through social cost reduction and approximate Nash equilibrium, respectively. Numerical experiments demonstrate the effectiveness of these approximations, offering insights for managing large-scale smart grid systems.
Appendix A Proofs of Auxiliary Lemmas
In this appendix, we collect the proofs of the auxiliary lemmas. We start with the proof of Lemma 2.1.
Proof of Lemma 2.1.
It follows from (2.16) that
$d\hat{X}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)\hat{X}_{t}^{i}+B(t)}{c}dt+\sigma_{i}dW_{t}^{i},\quad\hat{X}_{0}^{i}=x_{0}^{i},$
which implies that
$d\overline{\hat{X}}_{t}=-\frac{\bar{P}(t)+2a(t)\overline{\hat{X}}_{t}+B(t)}{c}dt+\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i},\quad\overline{\hat{X}}_{0}=\bar{x}_{0}^{N}.$
Gronwall Lemma, along with Assumption $\boldsymbol{(A_{2})}$, shows that there exists an $O(1)$ number $K$ such that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{X}_{t}^{i}|^{2}\right]\leq K,\ \quad\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\overline{\hat{X}}_{t}\right|^{2}\right]\leq K,\ \ \forall i=1,\ldots,N.$
Therefore, from (2.16), there exists an $O(1)$ number (still denoted by $K$) that satisfies
$\mathbb{E}\left[\int_{0}^{T}(\hat{v}_{t}^{i})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\hat{v}_{t}^{i})^{2}dt\right]\leq K,\ \forall i=1,\ldots,N,\$
$\mathbb{E}\left[\int_{0}^{T}\hat{Q}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}\hat{Q}_{t}^{2}\right]\leq K,\ \forall i=1,\ldots,N.$ (A.1)
For any $i$, we can consider all agents except for agent $i$ as a large population which consists of N1N-1 agents. Thus, by applying our previous calculations, we obtain
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N-1}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K.$
This leads to
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K.$
To proceed, we require an exact expression of $\hat{P}$. We directly solve from (2.2)
$\hat{P}_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha\hat{Q}_{s}ds.$
It follows that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]$ $\leq 2\mathbb{E}\left[\sup_{0\leq t\leq T}\left\{\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds\right)^{2}\right\}\right]$
$\quad+2\mathbb{E}\left[\sup_{0\leq t\leq T}\left\{\left(\int_{0}^{t}e^{-\alpha(t-s)}\alpha\hat{Q}_{s}ds\right)^{2}\right\}\right]$
$\leq O(1)+(\alpha T)^{2}\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{Q}_{t}}|^{2}\right]$
$\leq O(1),$
where we have applied (A). This shows that
$\mathbb{E}\left[\int_{0}^{T}\hat{P}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq O(1).$
Moreover, (A) gives
$\left|\mathbb{E}\left[\int_{0}^{T}\bigg{(}\frac{\eta}{2}(\hat{X}_{t}^{i}-\kappa)^{2}+\frac{c}{2}(\hat{v}_{t}^{i})^{2}\bigg{)}dt+\frac{\gamma}{2}(\hat{X}_{T}-\zeta)^{2}\right]\right|\leq O(1).$
The $O(1)$ bound for $\|\hat{v}\|$ and $\|\hat{P}\|$, together with Cauchy-Schwartz inequality, implies
$\left|\mathbb{E}\left[\int_{0}^{T}\hat{v}_{t}^{i}\hat{P}_{t}dt\right]\right|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}\|\leq O(1).$
This proves that
$|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq O(1),\quad i=1,\ldots,N.$
Let $K^{\prime}$ be an $O(1)$ number that exceeds $K$ and all the $O(1)$ numbers above, and we complete the proof of Lemma 2.1. ∎
Here are the proofs of Lemma 2.2 and Lemma 2.3.
Proof of Lemma 2.2.
Note that $v\in\mathcal{A}$ implies
$\int_{0}^{T}v_{t}^{2}dt<\infty,\ a.s.$
Consequently, we adopt $\int_{0}^{T}v_{t}^{2}dt<\infty$ in what follows, taking into account that events within a zero-probability set do not affect the expectation. In this setting, the integral as shown in Lemma 2.2 exists and is finite. To present the proof clearly, we start by introducing some notation. We set
$A_{1}=\begin{pmatrix}0&0\\ 0&{-\alpha}\end{pmatrix}\preceq O,\quad B_{1}=\begin{pmatrix}1\\ -\frac{\alpha}{N}\end{pmatrix},\quad C_{1}=\begin{pmatrix}{\epsilon}_{1}^{*}&1\end{pmatrix},\quad D_{1}={\epsilon}_{2}^{*},$
$x_{t}=\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix},\quad a^{*}={\epsilon}_{1}^{*}>0,\quad b^{*}=\frac{1}{2\alpha D_{1}}>0,\quad P_{1}=P_{1}^{\top}=diag{(a^{*},b^{*})}\succ O.$
Recalling from (2.18), we define
$y_{t}:=\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t}=C_{1}x_{t}+D_{1}v_{t}.$
Our objective is to apply Positive Real Lemma (Boyd et al. (1994)) to the system. To proceed, we denote the LMI for the system by
$L=\begin{pmatrix}A_{1}^{\top}P_{1}+P_{1}A_{1}&P_{1}B_{1}-C_{1}^{\top}\\ B_{1}^{\top}P_{1}-C_{1}&-D_{1}-D_{1}^{\top}\end{pmatrix}=\begin{pmatrix}0&0&a^{*}-\epsilon_{1}^{*}\\ 0&-2{\alpha b^{*}}&-\frac{b^{*}\alpha}{N}-1\\ a^{*}-\epsilon_{1}^{*}&-\frac{b^{*}\alpha}{N}-1&-2D_{1}\end{pmatrix}$.
We explore the conditions under which the LMI is semi-negative definite. By plugging in the exact value of $a^{*}$ and $b^{*}$, we discover that
$L\preceq O\iff O\preceq\begin{pmatrix}0&0&-a^{*}+\epsilon_{1}^{*}\\ 0&2{\alpha b^{*}}&\frac{b^{*}\alpha}{N}+1\\ -a^{*}+\epsilon_{1}^{*}&\frac{b^{*}\alpha}{N}+1&2D_{1}\end{pmatrix}=\begin{pmatrix}0&0&0\\ 0&\frac{1}{\epsilon_{2}^{*}}&\frac{1}{2\epsilon_{2}^{*}N}+1\\ 0&\frac{1}{2\epsilon_{2}^{*}N}+1&2\epsilon_{2}^{*}\end{pmatrix}$.
We conclude that
$L\preceq O\iff\text{det}\begin{pmatrix}\frac{1}{\epsilon_{2}^{*}}&\frac{1}{2\epsilon_{2}^{*}N}+1\\ \frac{1}{2\epsilon_{2}^{*}N}+1&2\epsilon_{2}^{*}\end{pmatrix}\geq 0\iff N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}.$
It is assumed that $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, so the LMI is semi-negative definite. Positive Real Lemma implies
$\int_{0}^{T}v_{t}(C_{1}x_{t}+D_{1}v_{t})dt\geq 0.$
By plugging in the exact form of $C_{1}$ and $D_{1}$ and taking the expectation, we immediately obtain that Lemma 2.2 holds true.
Proof of Lemma 2.3.
Fix positive numbers $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$, all independent of $N$, so that
$\epsilon_{2}=\frac{c}{8},\quad\epsilon_{3}<\frac{\eta}{4},\quad\epsilon_{4}+\epsilon_{5}<\frac{c}{8},\quad\epsilon_{3}\epsilon_{4}=\epsilon_{1}^{2}.$
Our first step relies on (2.2), which gives
$P_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\Big{(}\sum_{k\neq i}\hat{v}_{s}^{k}\Big{)}ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds.$
To evaluate how $P$ changes in response to $v$, we express $P$ as the sum of two components, one dependent on $v$ and the other independent of $v$. As shown below, we set
$P^{\prime}_{t}:=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\left(\sum_{k\neq i}\hat{v}_{s}^{k}\right)ds,$
$P_{t}^{\prime\prime}:=-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds.$
We naturally have $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$. The above expression for $P^{\prime\prime}_{t}$ is equivalently characterized by the dynamics $dP_{t}^{\prime\prime}=-\alpha P_{t}^{\prime\prime}-\frac{\alpha}{N}v_{t}^{i}$, $P^{\prime\prime}_{0}=0$. $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ are both processes induced by $v^{i}$, and the coefficients are similar to those in Lemma 2.2. However, to apply this lemma requires that the initial values for $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ equal zero, which is not possible to be fulfilled in the general case. Therefore, we need to take a detour. We construct the process below,
$d\dot{X}_{t}^{i}=v_{t}^{i}dt,\ \dot{X}_{0}^{i}=0.$
Note that the dynamics and initial value of $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ and that in Lemma 2.2 are identical. We then substitute $({X}_{t}^{*},\ P_{t}^{*})^{\top}$ with $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ in Lemma 2.2. It follows that, if $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$, we have
$\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}(\epsilon_{1}\dot{X}_{t}^{i}+P^{\prime\prime}_{t}+\epsilon_{2}v_{t}^{i})dt\right]\geq 0.$ (A.2)
Combining $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$ and (A.2), we obtain an inequality for $J_{i}$ satisfied by $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ $=\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}v_{t}^{i}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
$\geq\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}^{\prime\prime}v_{t}^{i}-|v_{t}^{i}P_{t}^{\prime}|\right)dt\right]$
$=\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}\right)dt\right]+\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}\left(\epsilon_{1}\dot{X}_{t}^{i}+P^{\prime\prime}_{t}+\epsilon_{2}v_{t}^{i}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{5}(v_{t}^{i})^{2}-|P_{t}^{\prime}v_{t}^{i}|\right)dt\right]$
$\geq\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{5}(v_{t}^{i})^{2}-|P_{t}^{\prime}v_{t}^{i}|\right)dt\right].$ (A.3)
From Lemma 2.1, direct calculation and the Cauchy-Schwartz inequality yield that
$\mathbb{E}\left[\sup_{t\in[0,T]}(P^{\prime}_{t})^{2}\right]$ $\leq 2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds\right)^{2}\right\}\right]$
$\quad+2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(\int_{0}^{t}\frac{\alpha}{N}e^{-\alpha(t-s)}\bigg{(}\sum_{k\neq i}\hat{v}_{s}^{k}\bigg{)}ds\right)^{2}\right\}\right]$
$\leq O(1)+(\alpha T)^{2}O(1)$
$\leq O(1).$
This implies that
$\mathbb{E}\left[\int_{0}^{T}(P^{\prime}_{t})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{t\in[0,T]}(P^{\prime}_{t})^{2}\right]\leq O(1).$
The fundamental inequality gives $|P^{\prime}_{t}v_{t}^{i}|\leq\frac{\epsilon_{5}}{2}(v_{t}^{i})^{2}+\frac{1}{2\epsilon_{5}}(P^{\prime}_{t})^{2}$, which leads to
$\mathbb{E}\left[\int_{0}^{T}\big{(}\epsilon_{5}(v_{t}^{i})^{2}-|P^{\prime}_{t}v_{t}^{i}|\big{)}dt\right]\geq\mathbb{E}\left[\int_{0}^{T}(\epsilon_{5}(v_{t}^{i})^{2}-\frac{1}{2}\epsilon_{5}(v_{t}^{i})^{2}-O(1)(P^{\prime}_{t})^{2})dt\right]\geq O(1).$
This simplifies (A) into the following inequality, which holds true once $N$ is larger than $\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq\mathbb{E}\left[\int_{0}^{T}\bigg{(}\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}\bigg{)}dt\right]+O(1).$
We want to cover the possible negative term $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}$ with a sum of several multiples of $(\dot{X}_{t}^{i})^{2}$ and $(v_{t}^{i})^{2}$. However, we only have $X_{t}^{i}$ in this inequality. Thus, we reduce $\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]$ into an expression of $\dot{X}_{t}^{i}$. We obtain from $(X_{t}^{i}-\kappa)-(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)=\dot{X}_{t}^{i}$,
$\mathbb{E}[(\dot{X}_{t}^{i})^{2}]$ $\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]+2\mathbb{E}[(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)^{2}]$
$\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)]^{2}+4(x_{0}^{i}-\kappa)^{2}+4\sigma_{i}^{2}T.$
This gives
$\mathbb{E}\left[(X_{t}^{i}-\kappa)^{2}\right]\geq$ $\frac{1}{2}\mathbb{E}\left[(\dot{X}_{t}^{i})^{2}\right]+O(1).$
Then,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ $\geq\mathbb{E}\left[\int_{0}^{T}\bigg{(}\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}\bigg{)}dt\right]+O(1)$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{(}\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg{)}dt$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\epsilon_{3}(X_{t}^{i}-\kappa)^{2}\right)dt\right]+O(1)$
$\geq\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{(}\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg{)}dt$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\right)dt\right]+O(1)\$
is satisfied for $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}=\frac{4(\sqrt{2}+1)}{c}$. From the restrictions on $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$, the contents of the first two parentheses provide the desired leading term. Observe that, by the restriction $\epsilon_{1}^{2}=\epsilon_{3}\epsilon_{4}$, fundamental inequality implies $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\geq 0$. Therefore, the proof is complete.
The proofs of Lemma 2.4 and Lemma 2.5 follow.
Proof of Lemma 2.4.
For $v\in\mathcal{A}$, we adopt the condition $\int_{0}^{T}v_{t}^{2}dt<\infty$ in the analysis below, as in the proof of Lemma 2.2. We introduce the notion
$A_{2}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}\preceq O,\quad B_{2}=\begin{pmatrix}1\\ -\alpha\end{pmatrix},\quad C_{2}=\begin{pmatrix}\alpha&1\end{pmatrix},\quad D_{2}=1.$
Recalling Lemma 2.4’s statement, we introduce the output function
$\tilde{y}_{t}:=v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}=C_{2}\tilde{x}_{t}+D_{2}.$
Then, the process simplifies to
$\frac{d}{dt}\tilde{x}_{t}=A_{2}\tilde{x}_{t}+B_{2},\quad\tilde{x}_{0}=(0,\ 0)^{\top}.$
We introduce the following notions
$\tilde{a}=\alpha,\quad\tilde{b}=\frac{1}{\alpha},\quad P_{2}=diag{(\tilde{a},\tilde{b})}\succ O,$
and set the LMI to be
$L_{2}=\begin{pmatrix}A_{2}^{\top}P_{2}+P_{2}A_{2}&P_{2}B_{2}-C_{2}^{\top}\\ B_{2}^{\top}P_{2}-C_{2}&-D_{2}-D_{2}^{\top}\end{pmatrix}=\begin{pmatrix}0&0&\tilde{a}-\alpha\\ 0&-2\alpha\tilde{b}&-\tilde{b}\alpha-1\\ \tilde{a}-\alpha&-\tilde{b}\alpha-1&-2\end{pmatrix}$.
We state that $L_{2}\preceq 0$, for it is equivalent to
$0\preceq\begin{pmatrix}0&0&-\tilde{a}+\alpha\\ 0&2\alpha\tilde{b}&\tilde{b}\alpha+1\\ -\tilde{a}+\alpha&\tilde{b}\alpha+1&2\end{pmatrix}=\begin{pmatrix}0&0&0\\ 0&2&2\\ 0&2&2\end{pmatrix}$.
This is obvious. From the Positive Real Lemma, we conclude that
$\int_{0}^{T}v_{t}\tilde{y}_{t}dt=\int_{0}^{T}v_{t}(v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t})dt\geq 0.$
We complete the proof by taking the expectation. ∎
Proof of Lemma 2.5.
Making use of Assumption $(\boldsymbol{A_{4}})$, we fix $\epsilon,\epsilon^{\prime}>0$ small enough, $A\in(0,\frac{\eta}{2})$, and $h\in(0,1)$ so that
$2\sqrt{\left(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon\right)Ah}>\alpha.$ (A.4)
In order to apply Lemma 2.4, which requires the starting value to be 0, we introduce three processes $P_{t}^{\dagger}$, $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$, whose dynamics satisfy
$d\overline{X}_{t}=Q_{t}dt+\frac{1}{N}\sum_{k=1}^{N}\sigma_{k}dW_{t}^{k}:=Q_{t}dt+\sigma dB_{t},\quad\overline{X}_{0}=\bar{x}_{0}^{N},\quad\bigg{(}\sigma=\frac{1}{N}\sqrt{\sum_{k=1}^{N}\sigma_{k}^{2}}\bigg{)},$
$d\overline{X}_{t}^{\prime}=Q_{t}dt,\quad\overline{X}_{0}^{\prime}=0,$
$d{P}_{t}^{\dagger}=\alpha(-Q_{t}-P_{t}^{\dagger})dt,\quad P_{0}^{\dagger}=0.$
Note that $\overline{X}_{t}$ is the average of $X_{t}^{i}$, so it contains a Brownian motion term and does not start at zero. The other two processes start at zero, and do not contain a Brownian motion term in their dynamics. From their definition, we get
$\overline{X}_{t}=\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime}+\sigma B_{t}.$
A simple calculation shows
$P_{t}^{\dagger}=-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds.$ (A.5)
The dynamics and initial value of $(\overline{X}_{t}^{\prime},\ P_{t}^{\dagger})^{\top}$ are the same as that of the process displayed in Lemma 2.4. Therefore, Lemma 2.4 implies that
$\mathbb{E}\left[\int_{0}^{T}Q_{t}(Q_{t}+P_{t}^{\dagger}+\alpha\overline{X_{t}}^{\prime})dt\right]\geq 0.$ (A.6)
Recall that
$J_{soc}(\boldsymbol{v})=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}v_{t}^{i}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right].$
(A.6) involves $P_{t}^{\dagger}$ and $\overline{X}_{t}^{\prime}$, which do not appear in $J_{soc}$. Therefore, we aim to reduce $J_{soc}$ into an expression where $\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ and part of $\frac{c}{2}v_{t}^{i}$ are substituted by expressions of $P_{t}^{\dagger}$, $Q_{t}$ and $\overline{X}_{t}^{\prime}$. Our first step relies on the convexity of $L$, which indicates that $\sum_{i=1}^{N}\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ is not smaller than $\frac{\eta N}{2}(\overline{X}_{t}-\kappa)^{2}$, and $\sum_{i=1}^{N}(\frac{c}{2}-\epsilon^{\prime})(v_{t}^{i})^{2}$ is not smaller than $\sum_{i=1}^{N}(\frac{c}{2}-\epsilon^{\prime})Q_{t}^{2}$. In light of this, we write
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\bigg{(}\frac{\eta}{2}(\overline{X}_{t})^{2}-\eta\kappa\overline{X}_{t}+\frac{\eta\kappa^{2}}{2}+\Big{(}\frac{c}{2}-\epsilon^{\prime}\Big{)}Q_{t}^{2}+P_{t}Q_{t}+\epsilon^{\prime}(v_{t}^{i})^{2}\bigg{)}dt\right].$ (A.7)
Our second step relies on the exact expression of $P$. To separate the control-independent components, we introduce the following definition
$C(t)=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds,\quad t\in[0,T].$
This function is obviously equipped with a $\|\cdot\|_{\infty}$ bound over $[0,T]$. The exact form of $P_{t}$, along with (A.5), shows that
$P_{t}$ $=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds$
$=P_{t}^{\dagger}+C(t).$ (A.8)
Our third step relies on the relationship between $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$. We have
$\mathbb{E}\left[(\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime})^{2}\right]$ $=\mathbb{E}\left[(\overline{X}_{t}-\sigma B_{t})^{2}\right]$
$\leq\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+\left(\frac{1}{h}-1\right)\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+\frac{h}{1-h}\mathbb{E}\left[\sigma^{2}B_{t}^{2}\right]+O(1)$
$=\frac{1}{h}\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+O(1).$
This implies that
$\mathbb{E}[(\overline{X}_{t})^{2}]\geq h\mathbb{E}[(\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime})^{2}]+O(1).$ (A.9)
Plug (A), (A.9) into (A.7), and rearrange the terms to get another inequality for $J_{soc}$
$J_{soc}(\boldsymbol{v})$ $\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\frac{\eta h}{2}\left({(\bar{x}_{0}^{N})}^{2}+2\overline{X_{t}}^{\prime}\bar{x}_{0}^{N}\right)-\eta\kappa\overline{X_{t}}^{\prime}+\left(\frac{\eta h}{2}-Ah\right)\overline{X_{t}}^{\prime 2}\right)dt\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X_{t}}^{\prime}+Ah\overline{X_{t}}^{\prime 2}+\left(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon\right)Q_{t}^{2}-\alpha Q_{t}\overline{X_{t}}^{\prime}\right)dt\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(C(t)Q_{t}+\epsilon Q_{t}^{2}+\epsilon^{\prime}{(v_{t}^{i})}^{2}\right)dt\right]+O(1).$
We have derived from Lemma 2.4 that $\mathbb{E}\left[\int_{0}^{T}(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X}_{t}^{\prime})dt\right]\geq 0$, and (A.4), together with the fundamental inequality, shows that $(Ah\overline{X}_{t}^{\prime 2}+(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon)Q_{t}^{2}-\alpha Q_{t}\overline{X}_{t}^{\prime})$ stays non-negative. These observations lead us to further reduce the inequality for $J_{soc}$ into
$J_{soc}(\boldsymbol{v})$ $\geq\frac{1}{N}\mathbb{E}\bigg{[}\sum_{i=1}^{N}\int_{0}^{T}\bigg{(}\frac{\eta h}{2}\Big{(}({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\Big{)}-\eta\kappa\overline{X}_{t}^{\prime}+\left(\frac{\eta h}{2}-Ah\right)\overline{X}_{t}^{\prime 2}$
$\qquad\qquad+\left(\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|\right)+\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\bigg{)}dt\bigg{]}+O(1).$
Note that $(\frac{\eta}{2}h-Ah)\overline{X}_{t}^{\prime 2}+\frac{\eta h}{2}\big{(}({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\big{)}-\eta\kappa\overline{X}_{t}^{\prime}$ and $\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|$ have an $O(1)$ lower bound, so
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\left(\int_{0}^{T}\left(\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\right)dt\right)\right]+O(1).$
The proof of Lemma 2.5 is complete.
Appendix B Proofs of Auxiliary Propositions
In this section, we collect the proofs of the propositions.
Proof of subsection 2.1.
First, we need to confirm that (2.10) admits a unique classical solution on $[0,T]$. To prove existence, let function
$Z(a)=-\frac{\eta}{2}+\frac{2a^{2}}{c},\quad a\in\mathbb{R}.$
Note that if $\gamma\neq\sqrt{c\eta}$, (2.10) is equivalent to
$\frac{da}{Z(a)}=dt,\quad a(T)=\frac{\gamma}{2}.$
By integrating the above equation from $T$ to $t$, and observing that if $\gamma=\sqrt{c\eta}$, the solution must be a constant, we obtain that
$\left\{\begin{aligned} &a(t)=-\frac{\sqrt{c\eta}}{2}\frac{1}{\tanh\left(\sqrt{\frac{\eta}{c}}(t-T)-\operatorname{arctanh}\left(\frac{\sqrt{c\eta}}{\gamma}\right)\right)},\quad\gamma>\sqrt{c\eta},\\ &a(t)\equiv\frac{\sqrt{c\eta}}{2},\quad\gamma=\sqrt{c\eta},\\ &a(t)=\frac{\sqrt{c\eta}}{2}\operatorname{tanh}\left(\sqrt{\frac{\eta}{c}}\left(t-T\right)+\operatorname{arctanh}\left(\frac{\gamma}{\sqrt{c\eta}}\right)\right),\quad\gamma<\sqrt{c\eta}.\end{aligned}\right$.
From the analysis above, we conclude that (2.10) gives a unique solution on $[0,T]$. In the following, it is convenient to explicitly compute from (2.11)
$B(t)=e^{-A(t)}\bigg{(}-e^{A(T)}\gamma\zeta-\int_{t}^{T}\left(\frac{2a(s)\bar{P}(s)}{c}+\eta\kappa\right)e^{A(s)}ds\bigg{)},$
where $A(t)$ is any primitive function for $\frac{2a(t)}{c}$, so (2.11) gives a unique solution on $[0,T]$. The solution to (2.12) obviously exists and is unique, once we integrate both sides from $T$ to $t$. The proof is complete. ∎
Proof of subsection 2.1.
Let
$\bar{k}(t)=\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix},\quad M(t)=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\end{pmatrix},\quad K=\begin{pmatrix}\eta\kappa\\ 0\\ \alpha\beta\end{pmatrix}$.
By Peano theorem, the initial condition gives a local solution. We need to verify that this solution extends to $t=T$ and does not blow up.
We introduce the following notation: the Frobenius norm of any matrix $L$ is denoted by $\|L\|_{\mathrm{F}}$. It is clear that $\|M(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$; we denote this bound by $M^{*}$, and denote $\|\bar{k}(t)\|_{\mathrm{F}}$ by $n(t)$. Multiply both sides of (2.13) with $2\bar{k}(t)^{\top}$, it follows that
$\frac{d}{dt}(n^{2}(t))=2(\bar{k}(t)^{\mathrm{T}}M(t)\bar{k}(t)+\bar{k}(t)^{\top}K)\leq 2n^{2}(t)M^{*}+n(t)\|K\|_{\mathrm{F}}.$
By basic ODE theorem, $n(t)$ stays below a function $N(t)$ characterized by
$d(N^{2}(t))=2(M^{*}N^{2}(t)+\|K\|_{\mathrm{F}}N(t)),\quad N(0)=n(0)+1.$
By a simple calculation, $N(t)$ has a uniform bound for $t\in[0,T]$, so $\bar{k}(t)$ stays finite on $[0,T]$. Therefore, the existence of the solution holds true. Moreover, $M(t)$ is Lipschitz continuous on $[0,T]$, so the uniqueness of the solutions naturally follows from the Picard-Lindelöf theorem. The proof is complete.
Proof of subsection 2.1.
We adopt the notation established in the proofs of subsection 2.1. Deducing in the same way as in the proof of subsection 2.1, we obtain the existence property. Recalling that $M(t)$ is Lipschitz continuous, we obtain the uniqueness property. ∎
Proof of subsection 2.1.
We adopt the notation established previously in subsection 2.1. By the definition of $\phi_{b_{0}}$, as shown in subsection 2.1 , we discover that for any $b_{0}\neq 0$, $\frac{\phi_{b_{0}}-\phi_{0}}{b_{0}}$ solves (2.15). Consequently, from the uniqueness property provided by subsection 2.1, we have
$\phi_{b_{0}}(t)-\phi_{0}(t)=b_{0}B_{1}(t).$
Let $t=T$. From $B_{1}(T)\neq 0$, we conclude that there exists a unique $b_{0}\in\mathbb{R}$ that satisfies $\phi_{b_{0}}(T)=-\zeta\gamma$. Thus, there exists only one solution to (2.13) and (2.14). The proof is complete. ∎
Proof of subsection 2.2.
Set
$G(t)=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}$.
Note that $G(t)$ is continuously differentiable for every $t\in[0,T]$, so the solution of each equation, if exists, must be unique. To prove existence, we observe that $\|G(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$, and work exactly as in subsection 2.1 to finish the proof. ∎
Proof of subsection 2.2.
Recalling from subsection 2.2 that $G(t)$ is continuously differentiable everywhere, we conclude that the solution, if exists, must be unique, by the Picard-Lindelöf theorem. The existence holds by the same way as in subsection 2.1. ∎
Proof of subsection 2.2.
We observe that , for $\forall b_{0},\ l_{0}\in\mathbb{R}$, $b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$ solves (2.28), and has boundary values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$. The uniqueness of solutions provided by subsection 2.2 indicates that $\phi_{b_{0},l_{0}}^{*}=b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$. Therefore, by letting $t=T$, $\phi_{b_{0},l_{0}}^{*}(T)=b_{0}\phi_{1}^{*}(T)+l_{0}\phi_{2}^{*}(T)+\phi_{0,0}^{*}(T)$. By Assumption $(\boldsymbol{A_{3}})$, there exists a unique pair of constants $b_{0},\ l_{0}\in\mathbb{R}$ such that the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a solution that satisfies (2.28) and (2.29). Thus, the solution for (2.28) and (2.29) exists and is uniquely determined. The proof is complete. ∎