### v2

Abstract
We study the optimal charging strategies for large-scale electric vehicles in smart grids within a finite-horizon mean field game framework. We first establish the existence and uniqueness of the solution to the consistency condition equation, which characterizes the optimal charging behavior in the mean field limit. Building on this result, we construct approximate optimal charging strategies for a finite population of vehicles in both non-cooperative and cooperative settings. Finally, we provide numerical analyses that illustrate and compare the approximate strategies in the non-cooperative and cooperative games.
Keywords: Smart grid, mean field game, mean field equilibrium, approximate Nash equilibrium.
MSC 2020: 91A06, 49L12, 49N80, 60H30
1 Introduction
As the global energy structure continues to shift towards low-carbonization, the number of electric vehicles (EVs) is growing rapidly. However, the spatiotemporal concentration of large-scale EV charging loads presents a significant challenge to the supply-demand balance of smart grids. In this context, developing a scientifically robust and highly efficient collaborative optimization strategy framework for EV charging is not only crucial for enhancing the operational resilience of the smart grid but also a core breakthrough for optimizing users’ electricity costs and achieving efficient allocation of energy resources (Sultan et al. (2022) and Yetkin et al. (2024)). Typically, the number of EVs requiring charging in the grid is enormous, leading to the “curse of dimensionality” when determining the optimal charging (or storage) strategy for each vehicle.
In the context of large-scale network structures, mean field games (MFGs) were independently proposed by Huang et al. (2006) and Lasry and Lions (2007), which devoted to the analysis of dynamic systems where a large number of players interact strategically with each other. By employing the “mean field” approximation, MFGs simplify the complex multi-player game problem into an interaction between a single player and the aggregate behavior of the population. This framework provides an effective analytical tool for addressing the interactions among players in large-scale systems. For more comprehensive insights into MFG theory, the reader may refer to Bensoussan et al. (2016), Carmona and Delarue (2018), Gomes and Saúde (2014) and the additional references cited within those works. Its application has spanned a wide range of fields including economics and finance  (Lacker and Zariphopoulou, 2019; Nuno and Moll, 2018), large networks or graphs (Lacker and Soret, 2022), operations research (Wang and Huang, 2019; Wang et al., 2022), engineering and machine learning (Zhou and Xu (2022)), epidemics control (Laguzet and Turinici (2015); Lee et al. (2021)), smart grid (Gomes and Saúde (2021)) and more.
Most of the MFG literature in smart grids focuses on two types of problems: Nash equilibria or social optimum. Nash equilibria in mean-field control has been investigated in Saldi et al. (2019) and Guo et al. (2022). The closest work to ours, authored by Cohen and Zell (2025), examines finite-state, infinite-horizon mean-field games with ergodic costs, demonstrates that solutions to the mean-field game system yield approximate Nash equilibria in corresponding finite-player games, and establishes a large deviation principle for empirical measures associated with these equilibria.
Social optimum solutions for MFG are also well documented. For example, Li et al. (2016) examine the relationship between mean-field games and social welfare optimization problems. Salhab et al. (2016) focus on dynamic collective choice problems, and address scenarios where a large number of players cooperatively choose between multiple destinations while influenced by group behavior. Feng et al. (2021) involve a major agent and numerous minor agents and investigates a mixed stochastic LQG social optimization.
Moreover, versatile framework for modeling price dynamics in markets with numerous interacting agents are applied in MFGs. Notable approaches are listed below: Gomes and Saúde (2021) introduce a price formation model where numerous small players can store and trade a commodity like electricity; Gomes et al. (2020) propose a MFG model for price formation of a commodity with production subject to random fluctuations. The dynamic game models have been extensively utilized to analyze duopolistic competition with sticky prices. Existing literature related to it includes Cellini and Lambertini (2004), Wiszniewska-Matyszkiel et al. (2015), Valentini et al. (2021), Kańska and Wiszniewska-Matyszkiel (2022) and Hoof (2021).
This paper aims to develop a strategic framework for a large market composed of numerous agents, using the methodology of MFG. Within a setting characterized by sticky prices and finite time horizons, we investigate how agents can optimize their behavior to minimize expected losses. In the cooperative game case, our objective is to reduce the average social cost across the market. On the other hand, under the non-cooperative framework, we focus on constructing an approximate Nash equilibrium that captures the decentralized decision-making of individual agents. This dual perspective allows for a comprehensive analysis of collective and competitive behaviors in large-scale smart grid systems.
Given the close resemblance between our model and the models of Gomes et al. (2023), we now emphasize the distinctions in both model structure and methodological approach. The model in Gomes et al. (2023) determines the mean process of the control variable through a stochastic differential equation. Subsequently, by employing variational methods, they seek a price process such that, when agents take optimal actions to minimize transaction costs, the market clears and supply meets demand. Unlike in their work, our price process is endogenously generated within the system and is a stochastic process driven by control variables. The control variables can be freely adjusted, and the agents’ commodities are subject to independent noise disturbances. The more fundamental difference lies in our objective: we aim to construct $\epsilon$-optimal solutions for both cooperative and non-cooperative games in a large system, rather than achieving supply-demand equilibrium. These distinctions lead us to adopt a completely different methodological approach in our analysis– our methodological approach mainly involves the standard MFE approach, that is, to approximate stochastic processes using fixed functions, rather than employing variational techniques.
Our mathematical contributions can be summarized in three aspects: First, within a finite-horizon framework, we propose a mean-field interaction cost function, through which we derive approximate solutions for both the cooperative and non-cooperative games, while the literature contains relatively few works that offer both approximate Nash equilibria and approximate social optima. Second, for specific matrix equations, we construct the solution induced by specific initial values and study the existence and uniqueness of the mean field through its properties. Most of the existing literature employs the Banach fixed-point theorem or other methods to find equilibrium points. In contrast to these approaches, our method is able to more precisely characterize the state of the mean field. Third, we apply the Positive Real Lemma to demonstrate that certain parameter restrictions can ensure that the coercivity condition holds, thereby effectively addressing the challenge posed by the positive correlation between the cost function and the product of price and demand term. To the best of our knowledge, this challenge has not been addressed in the existing literature.
The rest of the paper is organized as follows: In Section 2, we introduce the model of the smart grid, which consists of a large number of electric vehicles. In Section 3, we construct the $\epsilon$-Nash equilibrium charging strategies for the EVs in the non-cooperate games. Afterwards, in section 4, we construct the asymptotic social optimum charging strategies for the cooperate games. In Section 5, we give numerical simulations for the charging strategies proposed in both the non-cooperate games and cooperate games. Section 6 concludes.
2 The Model and the Optimal Strategies
In this section, we propose the sticky price model and the optimal charging strategies of smart grid with numerous agents (e.g. EVs). First, we introduce our model. Let the number of agents in the smart grid be $N$. The agents control their electricity (also referred to as commodities) through controlling the charging rate over a shared common time horizon $[0,T]$. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions, which supports the independent $N$-dimension standard Brownian motions $(W_{t}^{1},\cdots,W_{t}^{N})_{t\in[0,T]}$. Denote the electricity held by agent $i$ at time $t$ by $X_{t}^{i}$. Let the dynamics of $X^{i}$ satisfy
$dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i}\in\mathbb{R}.$ (2.1)
Here, the parameter $\sigma_{i}\geq 0$ measures the volatility of agent $i$’s electricity, and the stochastic noise term represents uncertainty or variability in the smart grid system, such as grid disturbances or physical noise, communication or control delays, or user behavior uncertainty. In addition, we assume that agent $i$ is equipped with an initial electricity $x_{0}^{i}\in\mathbb{R}$, As in (2.1), the control variable for agent $i$, also referred to as trading rate (or charging rate) is denoted by $v^{i}$. To introduce the admissible strategy set of $v^{i}$, we assign every stochastic control $w$ the norm $\left\|w\right\|=\sqrt{\mathbb{E}\left[\int_{0}^{T}w_{t}^{2}dt\right]}$. Then, $v^{i}$ should belong to the admissible strategy set
[…]
We then introduce the evolution of the price process $P$. We consider a dynamic oligopoly setting, where prices do not instantly adjust to changes in market conditions, such as shifts in supply or demand. Instead, its evolution is affected by its current price and average trading rate. A natural way of modeling the sticky price in dynamic oligopoly is (Cellini and Lambertini (2004) and Wiszniewska-Matyszkiel et al. (2015))
$dP_{t}=\alpha(\beta-Q_{t}-P_{t})dt,\quad P_{0}=p_{0},$ (2.2)
where $\alpha,\beta>0$ are constant market factors, and the initial price is $p_{0}$. The parameter $\alpha$ measures the sensibility of $P$ to current situations, and $Q_{t}:=\frac{1}{N}\sum_{i=1}^{N}v_{t}^{i}$ is the average trading rate at time $t$. The sticky price model well captures the persistent effects of the past price conditions on the current prices, reflecting realistic frictions in electricity markets.
In the following, by referring to existing literature, we propose a reasonable cost function for each agent in the game. Inspired by Gomes et al. (2023), we let the expected overall cost of agent $i$ be of the following form
$J_{i}(v^{i},\boldsymbol{v}^{-i})=\mathbb{E}\left[\int_{0}^{T}\left(L(X_{t}^{i},v_{t}^{i})+P_{t}v_{t}^{i}\right)dt+\Psi(X_{T}^{i})\right],$ (2.3)
where $\boldsymbol{v}^{-i}:=\{v^{1},\cdots,v^{i-1},v^{i+1},\cdots,v^{N}\}$. We assume the running cost and the terminal cost to be
$L(x,v)=\frac{\eta}{2}(x-\kappa)^{2}+\frac{c}{2}v^{2}\quad\mathrm{and}\quad\Psi(x)=\frac{\gamma}{2}(x-\zeta)^{2}$
respectively. The parameter $\zeta$ corresponds to the preferred final storage, and $\kappa$ is the preferred instantaneous storage. We make the assumption that $\eta>0,\ \gamma\geq 0,\ c>0,\ \kappa\in\mathbb{R},\ \text{and}\ \zeta\in\mathbb{R}$ are constants. Compared with the conventional LQG framework, this cost function includes an additional $vP$ term. The introduction of this term weakens the coercivity condition, and increases the difficulty of estimating $\epsilon_{N}$ and deriving the optimal strategies.
Finally, building on the previously proposed cost function, we introduce an objective function for the system. We consider two different game cases. In one case (namely, the non-cooperate game case), each agent controls its commodity independently, and does not take into account how its actions might affect others’ expected cost. In other words, agent $i$ only aims to minimize its expected value function $J_{i}$, and does not consider how $J_{k}(\forall 1\leq k\leq N,k\neq i)$ changes according to its actions. Under this setting, it is natural for us to find an $\epsilon$-Nash equilibrium for the system.
A Nash equilibrium is a key concept in game theory that represents a stable state in a strategic interaction where no player can improve their outcome by unilaterally changing their strategies. It occurs when each player’s strategy is optimal given the strategies chosen by all other players. However, as $N$ enlarges, it becomes rather challenging to find the exact Nash equilibrium solution, so we aim to find a set of strategies that approaches the best outcome as $N$ enlarges, referred to as a set of $\epsilon$-Nash equilibrium strategies. An $\epsilon$-Nash equilibrium characterizes a situation where each player’s chosen strategy is almost the best response to the strategies of the others, with the difference between the expected personal cost of this strategy and that under the optimal strategy not exceeding a positive number $\epsilon_{N}$, which vanishes as $N$ goes to $\infty$. We introduce the precise definition of an $\epsilon$-Nash equilibrium:
Definition 2.1.
A set of admissive strategies $\hat{\boldsymbol{v}}=\{\hat{v}^{1},\cdots,\hat{v}^{N}\}\in\Pi_{i=1}^{N}\mathcal{A}^{i}$ is an $\epsilon$-Nash equilibrium if
$J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})-\epsilon_{N}\leq\inf_{v^{i}\in\mathcal{A}_{i}}J(v^{i},\hat{\boldsymbol{v}}^{-i})\leq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\quad\forall i=1,\ldots,N,$ (2.4)
where $\epsilon_{N}$ goes to 0 as $N\to\infty$.
Under the second case (namely, the cooperate game case), instead of controlling the electricity independently, the agents cooperate with each other, and aim to achieve the best collective outcome. In this article, the collective outcome represents the average value of $J_{i}(v^{i},\boldsymbol{v}^{-i})$. It reads
$J_{soc}(v^{1},v^{2},\cdots,v^{N})=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(L(X_{t}^{i},v_{t}^{i})+P_{t}v_{t}^{i}\right)dt+\Psi(X_{T}^{i})\right].$
In large-scale systems, finding the exact social optimum becomes complex, so we focus on finding a set of strategies that approaches the best collective outcome as the number of agents tends to infinity. More specifically, we introduce the following definition:
Definition 2.2.
We say that a set of admissive strategies $\{\check{v}^{i}\}_{i=1}^{N}$ is a set of $\epsilon$-optimal social strategies if
$|J_{soc}(\check{v}^{1},\check{v}^{2},\cdots,\check{v}^{N})-\inf_{v^{i}\in\mathcal{A}_{i},1\leq i\leq N}J_{soc}(v^{1},v^{2},\cdots,v^{N})|\leq\epsilon_{N},$ (2.5)
where $\epsilon_{N}$ goes to 0 as $N\to\infty$.
Here, the restriction for $\epsilon_{N}$ reflects that the expected social cost under the proposed strategies converges to that of the social optimal strategies as $N$ goes to $\infty$.
It is worth emphasizing that asymptotic social optimum strategies may not be $\epsilon$-Nash equilibrium strategies. Under the asymptotic social optimum strategies, an agent may significantly reduce its own expected cost by only changing its own strategy, while causing even more cost for others, and comprehensively, give rise to the overall expected cost.
3 Non-cooperate Game
In this section, we aim to find an $\epsilon$-Nash equilibrium for this system. We will use a mean-field approximation approach, in which we construct $C[0,T]$ functions to approximate certain processes. Our approach can be outlined in the following steps.
  1. 1.
    The first step in constructing a mean-field approximation is to approximate $Q$ with a continuous function $\bar{Q}=(\bar{Q}(t))_{t\in[0,T]}$. It follows from (2.2) that the approximate price function $(\bar{P}(t))_{t\in[0,T]}$ satisfies
    $d\bar{P}(t)=\alpha(\beta-\bar{Q}(t)-\bar{P}(t))dt,\quad\bar{P}(0)=p_{0}.$ (3.1)
  2. 2.
    Replace $P_{t}$ by $\bar{P}(t)$ in objective function $J_{i}$ to obtain an auxiliary objective function
    $\bar{J}_{i}(v^{i};\bar{P})=\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right],$ (3.2)
    subject to state process $(X_{t})_{t\in[0,T]}$, which satisfies (2.1).
  3. 3.
    For each agent $i$, solve the auxiliary control problem (3.2) to find the optimal control $\hat{v}^{i}$. We calculate the average of the corresponding optimal controls and remove the noise term to obtain a function which measures the expectation of the average of the optimal controls. By requiring that this function equal $\bar{Q}$, we determine the exact form of $\bar{Q}$.
  4. 4.
    Prove that under certain assumptions, the set of controls derived from $\bar{Q}$ is indeed an $\epsilon$-Nash equilibrium.
Remark 1.
The rationale for proposing the first two steps is primarily based on the following reasons. Firstly, when the population enlarges to $\infty$, the representative agent has no influence on $Q$, and therefore has no influence on $P$, as but one agent amid a continuum. From its perspective, the process $P$ should be evolved in its cost as a deterministic function. Secondly, the optimization problem for (2.3) becomes much easier to solve once we treat $P$ as a fixed continuous function, instead of a stochastic process. For each $\bar{Q}$, a unique process $\bar{P}$ is determined by (3.1). Each agent determines its best response to minimize (3.2). By calculating the expectation of the average of the optimal controls, we derive a function determined by $\bar{Q}$. If the expectation of this function is equal to $\bar{Q}$, a closed-loop is formed. Within this loop, the strategy adopted by each agent yields outcomes that closely approximate those of the actual optimal strategy as the number of agents increases.
For the given function $\bar{Q}\in C[0,T]$, we derive the approximate price process $\bar{P}$ by (3.1). We propose the value function of the auxiliary control problem (3.2),
$K_{i}(t,x)=\inf_{v\in\mathcal{B}_{i}}\mathbb{E}\left[\int_{t}^{T}(L(X_{s},v_{s})+\bar{P}(s)v_{s})ds+\Psi(X_{T})\right].$
where $t$ represents the starting time, $X$ satisfies
$dX_{s}=v_{s}ds+\sigma_{i}dW_{s},\quad\forall t\leq s\leq T;\quad X_{t}=x,$ (3.3)
and the admissible set for $v$ is denoted by
[…]
By dynamic programming principle, the HJB equation satisfies
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{B}_{i}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$ (3.4)
along with the terminal condition $K_{i}(T,x)=\frac{\gamma}{2}(x-\zeta)^{2}$. We make the ansatz ${K}_{i}(t,x)=a_{i}(t)x^{2}+B_{i}(t)x+F_{i}(t)$ yields. Furthermore, the optimal feedback control achieving the minimum is given by
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a_{i}(t)x+B_{i}(t)}{c},\quad\forall t\in[0,T].$ (3.5)
By substituting (3.5) into (3.4), we obtain that
$\qquad\qquad\qquad a^{\prime}_{i}(t)+\frac{\eta}{2}-\frac{(2a_{i}(t))^{2}}{2c}=0,$ $a_{i}(T)=\frac{\gamma}{2},$ (3.6)
$\qquad\qquad\qquad B^{\prime}_{i}(t)-\eta\kappa-\frac{2a_{i}(t)(\bar{P}(t)+B_{i}(t))}{c}=0,$ $B_{i}(T)=-\gamma\zeta,$ (3.7)
$\qquad\qquad\qquad F^{\prime}_{i}(t)-\frac{(\bar{P}(t)+B_{i}(t))^{2}}{2c}+\frac{\eta\kappa^{2}}{2}+\sigma_{i}^{2}a(t)=0,$ $F_{i}(T)=\frac{\gamma\zeta^{2}}{2}.$ (3.8)
We will subsequently verify that (3.6)–(3.8) indeed admit a solution. To this end, we introduce section 3.
Proposition 3.1.
The solution to (3.6), (3.7) and (3.8) exists and is unique.
Remark 2.
Throughout this paper, unless stated otherwise, we seek classical solutions to the equations under consideration. Since the solution to (3.6) and (3.7) is uniform over $i$, so in the following, we abbreviate $a_{i}(t),B_{i}(t)$ as $a(t),B(t)$, respectively.
The proofs of the Propositions in this paper are collected to Appendix B.
Next, we present the verification theorem for the best-response control that minimizes (3.2).
Theorem 3.2.
(Verification theorem) We introduce the feedback control function as follows:
$\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t){X}_{t}^{i}+B(t)}{c},\quad\forall t\in[0,T],$
where $a(t),B(t)$ are given by (3.6) and (3.7). Then, the SDE
$d{X}_{t}^{i}=\hat{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad{X}_{0}^{i}=x_{0}^{i}$
admits a unique strong solution, denoted by $\hat{X}_{t}^{i}$, and $\hat{v}^{i}$ is an optimal Markovian control for minimizing (3.2).
Proof.
$\hat{v}:(t,x)\to-\frac{\bar{P}(t)+2a(t)x+B(t)}{c}$ is Lipschitz-continuous w.r.t $(t,x)\in[0,T]\times\mathbb{R}$, implying that (2.1) admits a unique strong solution (Theorem 8.3 in Le Gall, J. F. (2016)). Since $a(t)$, $B(t)$ and $F_{i}(t)$ are bounded on $[0,T]$, the function $K_{i}(t,x)$ satisfies the quadric growth condition; that is, there exists $C>0$ such that $|K_{i}(t,x)|\leq C(1+x^{2}),\ \forall(t,x)\in[0,T]\times\mathbb{R}$. Additionally, the terminal condition $K_{i}(T,\cdot)=\Psi(\cdot)$ is fulfilled. Under these conditions on $K_{i}$, together with
$\frac{\partial{K}_{i}(t,x)}{\partial t}+\inf_{v^{i}\in\mathcal{A}_{i}}\left\{v_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({v_{t}^{i}})^{2}+v_{t}^{i}\bar{P}(t)\right\}+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}$
$=\frac{\partial{K}_{i}(t,x)}{\partial t}+\hat{v}_{t}^{i}\frac{\partial{K}_{i}(t,x)}{\partial x}+\frac{c}{2}({\hat{v}_{t}^{i}})^{2}+\hat{v}_{t}^{i}\bar{P}(t)+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}{K}_{i}(t,x)}{\partial x^{2}}+\frac{\eta}{2}(x-\kappa)^{2}=0,$
we obtain the result by applying the arguments in Theorem 3.5.2, Pham (2009). ∎
We now approximate the average commodity process by a deterministic function. The rationale behind this approximation is as follows. Let $\overline{X}_{t}$ denote the average of $X_{t}^{i}$. Then the average optimal control takes the form
$\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c},$
which implies that the mean trading process evolves as follows
$d\overline{X}_{t}=-\frac{2a(t)\overline{X}_{t}+{B}(t)+\bar{P}(t)}{c}dt+\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i},\quad\overline{X}_{0}=\overline{x}_{0}^{N}.$
Note that $\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i}$ represents a Brownian motion with magnitude $\frac{1}{N}\sqrt{\sum_{i=1}^{N}\sigma_{i}^{2}}$. This magnitude clearly vanishes to zero as $N\to\infty$ provided that $\sigma_{i}\ (\forall i=1,2,\ldots,N)$ are bounded by a constant independent of $N$. Hence, under the standard mean-field approximation, we replace the average stochastic process by a deterministic function $\bar{x}(t)$, whose dynamics satisfy
$d\bar{x}(t)=-\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N}.$
By consistency condition, the expectation of the induced average trading process must satisfy
$\frac{2a(t)\bar{x}(t)+B(t)+\bar{P}(t)}{c}=\bar{Q}_{t},\quad\forall t\in[0,T].$
Using the relationship between $\bar{P}(t),B(t)$ and $\bar{Q}(t)$, this requirement is equivalent to the following system:
$\left\{\begin{aligned} &d{\bar{P}(t)}=\alpha(\beta-\bar{P}(t)-\bar{Q}(t))dt,\quad\bar{P}(0)=p_{0},\\ &B^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{P}(t)+B(t))}{c}=0,\quad B(T)=-\gamma\zeta,\\ &d\bar{x}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N},\\ &\bar{Q}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}.\end{aligned}\right$.
By introducing matrix notation, we obtain an alternative form, which requires the following function system
$\frac{d}{dt}\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix}=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\par\par\end{pmatrix}\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix}+\begin{pmatrix}\eta\kappa\\ 0\\ \alpha\beta\par\end{pmatrix}$, (3.9)
to hold simultaneously with the following initial or terminal conditions
$\begin{pmatrix}B(T),\ \bar{x}(0),\ \bar{P}(0)\end{pmatrix}^{\top}=\begin{pmatrix}-\zeta\gamma,\ \bar{x}_{0}^{N},\ p_{0}\end{pmatrix}^{\top}$. (3.10)
The existence of solutions to (3.9) and (3.10) is essential, as it determines whether the consistency requirement in Step 3 can be fulfilled. The main challenge here is: the boundary conditions are imposed at different time points.
To tackle this issue, we first study solutions with initial conditions specified at the same time point, and establish their properties in section 3 and section 3. These results serve as preparatory steps for section 3, where we prove that, under Assumption $\boldsymbol{(A_{1})}$ stated below, the coupled system (3.9)–(3.10) admits a unique solution.
Proposition 3.3.
[…]
section 3 shows that once the conditions $\bar{x}(0)=\bar{x}_{0}^{N}$ and $\bar{P}(0)=p_{0}$ are given, the solution to (3.9) is uniquely determined by the choice of $B(0)=b_{0}$ . Therefore, our remaining task is to adjust the value of $b_{0}$ so that $B(T)=-\gamma\zeta$ holds.
To this end, we construct the following auxiliary function system,
$\frac{d}{dt}\begin{pmatrix}B_{1}(t)\\ \bar{x}_{1}(t)\\ \bar{P}_{1}(t)\end{pmatrix}=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}B_{1}(t)\\ \bar{x}_{1}(t)\\ \bar{P}_{1}(t)\end{pmatrix},\forall t\in[0,T];\begin{pmatrix}B_{1}(0)\\ \bar{x}_{1}(0)\\ \bar{P}_{1}(0)\end{pmatrix}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$. (3.11)
The existence and uniqueness of (3.11) are established below in section 3.
Proposition 3.4.
There exists a unique solution to (3.11).
We now introduce Assumption $\boldsymbol{(A_{1})}$ and show that, under this assumption, the coupled system (3.9)–(3.10) indeed admits a solution.
Assumption $\boldsymbol{(A_{1}):}$ The solution to (3.11) satisfies $B_{1}(T)\neq 0$.
Proposition 3.5.
If Assumption $\boldsymbol{(A_{1})}$ holds, then (3.9), combined with (3.10), gives a unique solution.
We further impose an assumption that constrains the magnitudes of the elements in the sets $\{\sigma_{i}\}_{1\leq i\leq N}$ and $\{x_{0}^{i}\}_{1\leq i\leq N}$.
Assumption $\boldsymbol{(A_{2}):}$ The sets $\{x_{0}^{i}\}_{i=1}^{N}$ and $\{\sigma_{i}\}_{i=1}^{N}$ are respectively defined on two fixed compact sets that are independent of $N$.
Now, we propose the main theorem for this section.
Theorem 3.6.
Assume that Assumption $\boldsymbol{(A_{1})}$ and $\boldsymbol{(A_{2})}$ hold, and the unique solution to (3.9) and (3.10) is given by (B(t),x¯(t),P¯(t))(B(t),\ $\bar${x}(t),\ $\bar${P}($)^{\top}, t$hen the following system
$\left\{\begin{aligned} &d\hat{P}_{t}=\alpha(\beta-\hat{Q}_{t}-\hat{P}_{t})dt,\quad\hat{P}_{0}=p_{0},\\ &\hat{v}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)\hat{X}_{t}^{i}+B(t)}{c},\\ &d\hat{X}_{t}^{i}=\hat{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad\hat{X}_{0}^{i}=x_{0}^{i},\\ &\hat{Q}_{t}=\frac{1}{N}\sum_{i=1}^{N}\hat{v}_{t}^{i}.\end{aligned}\right$. (3.12)
is an $\epsilon$-Nash equilibrium. Moreover, specific calculations show that under the proposed strategies $\{\hat{v}_{t}^{i}\}_{i=1}^{N}$, $\epsilon_{N}=O\Big(\frac{1}{\sqrt{N}}\Big)$ in (2.4).
To prove Theorem 3.6, we will need three auxiliary lemmas. As a first step, in order to derive uniform bounds for the norms of the state processes under the optimal control and for the quantities $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),\ \forall i=1,2,\ldots,N$, we introduce Lemma 3.1.
Lemma 3.1.
There exists a positive constant $K^{\prime}$, independent of $N$ and $i$, such that
$\left\{\begin{aligned} &\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{X}_{t}^{i}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\int_{0}^{T}(\hat{v}_{t}^{i})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\hat{v}_{t}^{i})^{2}dt\right]\leq K^{\prime},\quad i=1,2,\ldots,N,\\ &\mathbb{E}\left[\sup_{0\leq t\leq T}|\overline{\hat{X_{t}}}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K^{\prime}\ ,\quad i=1,2,\ldots,N,\\ &\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq K^{\prime},\ \quad\mathbb{E}\left[\int_{0}^{T}\hat{Q}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{Q}_{t}|^{2}\right]\leq K^{\prime},\\ &|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq K^{\prime},\quad i=1,2,\ldots,N,\end{aligned}\right$.
where $\overline{\hat{X}}_{t}$ is the average of $\hat{X}_{t}^{i}$, from $i=1$ to $N$.
The proofs of all lemmas have been left to Appendix A.
[…]
$dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i},$
and the price dynamics become
$d{P}_{t}=\alpha\bigg(\beta-\frac{v_{t}^{i}}{N}-\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}-{P}_{t}\bigg)dt,\quad{P}_{0}=p_{0}.$
We aim to compare $J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$ with $J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ and establish (2.4) with $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$, or equivalently,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg(\frac{1}{\sqrt{N}}\bigg),\quad\forall\ v^{i}\in\mathcal{A}_{i}.$ (3.13)
To establish (3.13), we require a coercivity condition, whose precise formulation is given in Lemma 3.3. Before presenting it, we first introduce Lemma 3.2, which will streamline the proof of the coercivity result.
Lemma 3.2.
For any $N\in\mathbb{N}^{+}$ and $v\in\mathcal{A}_{i}$, we construct the process
$\frac{d}{dt}\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix}=\begin{pmatrix}0&0\\ 0&{-\alpha}\end{pmatrix}\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix}+\begin{pmatrix}1\\ -\frac{\alpha}{N}\end{pmatrix}v_{t},\ \quad X_{0}^{*}=0,\ P_{0}^{*}=0.$
Then, for any given $\epsilon_{1}^{*}>0$, $\epsilon_{2}^{*}>0$ and $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, we have
$\mathbb{E}\left[\int_{0}^{T}v_{t}(\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t})dt\right]\geq 0.$ (3.14)
Lemma 3.3.
There exists a constant $C_{0}$, which is independent of $N$ and $i$, such that the coercive condition
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq\mathbb{E}\left[\int_{0}^{T}\Big(\frac{\eta}{4}(X_{t}^{i}-\kappa)^{2}+\frac{c}{4}(v_{t}^{i})^{2}\Big)dt\right]+C_{0}$
is satisfied for any $N>\frac{4(\sqrt{2}+1)}{c}$.
We now return to our proof of Theorem 3.6.
Proof of Theorem 3.6.
According to Lemma 3.1, we get
$|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq K^{\prime}.$
For all $N>\frac{4(\sqrt{2}+1)}{c}$, if $\frac{c}{4}\|v^{i}\|^{2}+C_{0}\geq K^{\prime}$, then by Lemma 3.3, we immediately obtain that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i}),$
which yields (3.13). Otherwise, $\|v^{i}\|^{2}$ is bounded by $\frac{4(K^{\prime}-C_{0})}{c}$. Hence, it suffices to establish (3.13) in the case where $\|v^{i}\|$ is bounded by an $O(1)$ constant.
As noted earlier, $\bar{P}(t)$ is constructed for approximating $P_{t}$ and $\hat{P}_{t}$. This motivates us to express the difference of the expected cost as
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})$
$=\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right]-\mathbb{E}\left[\int_{0}^{T}(L(\hat{X}_{t}^{i},\hat{v}_{t}^{i})+\bar{P}(t)\hat{v}_{t}^{i})dt+\Psi(\hat{X}_{T}^{i})\right]$
$\quad+\mathbb{E}\Big[\int_{0}^{T}v_{t}^{i}(P_{t}-\bar{P}(t))dt\Big]-\mathbb{E}\Big[\int_{0}^{T}\hat{v}_{t}^{i}(\hat{P}_{t}-\bar{P}(t))dt\Big]$
$:=I_{1}+I_{2}-I_{3},$
where
$I_{1}$ $=\mathbb{E}\left[\int_{0}^{T}(L({X}_{t}^{i},{v}_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi({X}_{T}^{i})-\left(\int_{0}^{T}(L(\hat{X}_{t}^{i},\hat{v}_{t}^{i})+\bar{P}(t)\hat{v}_{t}^{i})dt+\Psi(\hat{X}_{T}^{i})\right)\right],$
$I_{2}$ $=\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}(P_{t}-\bar{P}(t))dt\right],$
$I_{3}$ $=\mathbb{E}\left[\int_{0}^{T}\hat{v}_{t}^{i}(\hat{P}_{t}-\bar{P}(t))dt\right].$
Thus, to finish the proof, it suffices to show
$I_{1}\geq 0,\quad I_{2}=O\Big(\frac{1}{\sqrt{N}}\Big),\quad I_{3}=O\Big(\frac{1}{\sqrt{N}}\Big).$ (3.15)
Since we have already established that $\hat{v}_{t}^{i}$ is optimal for minimizing
$\mathbb{E}\left[\int_{0}^{T}(L(X_{t}^{i},v_{t}^{i})+\bar{P}(t)v_{t}^{i})dt+\Psi(X_{T}^{i})\right],$
it follows directly that $I_{1}\geq 0$.
We now address the remaining term in (3.15). Direct calculation yields
$d\overline{\hat{X}}_{t}=-\frac{\bar{P}(t)+2a(t)\overline{{\hat{X}}}_{t}+B(t)}{c}dt+\frac{1}{N}\sum_{k=1}^{N}\sigma_{k}dW_{t}^{k},\quad\overline{\hat{X}}_{0}=\bar{x}_{0}^{N},$
$d\bar{x}(t)=-\frac{\bar{P}(t)+2a(t)\bar{x}(t)+B(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N},$
$\hat{P}_{t}-\bar{P}(t)=\frac{2\alpha}{c}\int_{0}^{t}e^{-\alpha(t-s)}a(s)\big(\overline{\hat{X_{s}}}-\bar{x}(s)\big)ds.$
Applying Grönwall’s lemma and Assumption $\boldsymbol{(A_{2})}$ yield
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\bar{x}(t)-\overline{\hat{X}}_{t}\right|^{2}\right]\leq\mathbb{E}\left[O(1)\frac{1}{N^{2}}\left(\sqrt{\sum_{k=1}^{N}\sigma_{k}^{2}}\right)^{2}\right]=O\left(\frac{1}{N}\right),$
which further implies, using the Cauchy–Schwarz inequality,
$\|\hat{P}-\bar{P}\|^{2}$ $\leq O(1)\mathbb{E}\left[\int_{0}^{T}\bigg(\int_{0}^{t}\left(e^{-\alpha(t-s)}\right)^{2}ds\bigg)\bigg(\int_{0}^{t}\left(\overline{\hat{X}}_{s}-\bar{x}(s)\right)^{2}ds\bigg)dt\right]$
$\leq O(1)\left(\int_{0}^{T}\bigg(\int_{0}^{t}\left(e^{-\alpha(t-s)}\right)^{2}ds\bigg)dt\right)\mathbb{E}\left[\int_{0}^{T}\left(\overline{\hat{X}}_{s}-\bar{x}(s)\right)^{2}ds\right]$
$=O\left(\frac{1}{N}\right).$
Then, from (2.2),
$P_{t}-\hat{P}_{t}=-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}(v_{s}^{i}-\hat{v}_{s}^{i})ds.$
From the price process (2.2), we deduce by another application of the Cauchy–Schwarz inequality that
$\|P-\hat{P}\|^{2}\leq\frac{1}{N^{2}}\left(\int_{0}^{T}dt\int_{0}^{t}(e^{-\alpha(t-s)}\alpha)^{2}ds\right)\mathbb{E}\left[\int_{0}^{T}(v_{s}^{i}-\hat{v}_{s}^{i})^{2}ds\right]=O\bigg(\frac{1}{N^{2}}\bigg),$
where the last equality follows from the fact that $\|\hat{v}^{i}\|$ and $\|v^{i}\|$ are all bounded by $O(1)$ numbers. These estimates imply
$\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),\quad\|P-\hat{P}\|=O\bigg(\frac{1}{N}\bigg),\quad\|P-\bar{P}\|\leq\|P-\hat{P}\|+\|\hat{P}-\bar{P}\|=O\left(\frac{1}{\sqrt{N}}\right),$
where the last step applies the triangle inequality. Applying the Cauchy-Schwartz inequality once more gives
$|I_{2}|\leq\|v^{i}\|\cdot\|P-\bar{P}\|=O\bigg(\frac{1}{\sqrt{N}}\bigg),$
$|I_{3}|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}-\bar{P}\|=O\bigg(\frac{1}{\sqrt{N}}\bigg),$
which establishes (3.15). Finally, combining all the above arguments, we conclude that
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})-J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})\geq O\bigg(\frac{1}{\sqrt{N}}\bigg).$
Hence, the proof is complete. ∎
4 The Cooperate Games
In this section, we aim to identify a set of strategies that achieve an $\epsilon$-optimal expected social cost. Since our objective is to minimize $J_{soc}$, it is essential to quantify the influence of agent $i$’s control on the value of $J_{soc}$. To this end, we decompose the price process $P$ into components that are affected and unaffected by agent $i$, and then analyze how $v^{i}$ contribute to $J_{soc}$.
Our approach can be outlined in the following steps.
  1. 1.
    We first show that under the optimal strategies of $J_{soc}$, the control of each agent $i$’ must minimize the following auxiliary control problem:
    $\check{J}_{i}(v^{i})$ $:=\mathbb{E}\left[\int_{0}^{T}\bigg(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}P_{t}-\frac{u_{t}^{i}}{N}\sum_{k=1,k\neq i}^{N}\check{v}_{t}^{k}\bigg)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right],$ (4.1)
    subject to
    $\left\{\begin{aligned} &\frac{dP_{t}}{dt}=\alpha\bigg(\beta-\frac{1}{N}\sum_{k\neq i}^{N}\check{v}_{t}^{k}-\frac{1}{N}v_{t}^{i}-P_{t}\bigg),\quad P_{0}=p_{0},\\ &dX_{t}^{i}=v_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad X_{0}^{i}=x_{0}^{i},\\ &u_{t}^{i}=\int_{0}^{t}e^{-\alpha(t-s)}\alpha v_{s}^{i}ds.\end{aligned}\right$. (4.2)
  2. 2.
    We approximate the empirical mean $\frac{1}{N}\sum_{k\neq i}\check{v}_{t}^{k}$ with a continuous function $(\bar{q}(t))_{0\leq t\leq T}$. Subsequently, we construct a continuous approximation function $(\bar{p}(t))_{0\leq t\leq T}$ to the price process $(P_{t})_{t\in[0,T]}$, which evolves according to
    $d\bar{p}(t)=\alpha(\beta-\bar{p}(t)-\bar{q}(t))dt,\quad\bar{p}(0)=p_{0}.$ (4.3)
  3. 3.
    For each agent $i=1,\ldots,N$, applying the above approximation yields the following auxiliary objective function,
    $J^{\prime}_{i}(v^{i};\bar{p},\bar{q})=\mathbb{E}\left[\int_{0}^{T}\Big(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)-\bar{q}(t)u_{t}^{i}\Big)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right],$ (4.4)
    subject to the state dynamics $(X^{i},u^{i})$ given by (4.2).
  4. 4.
    By solving the control problem $J^{\prime}_{i}$, we obtain the corresponding optimal strategies and derive the consistency condition that the function $\bar{q}$ must satisfy. Finally, we verify that the strategies generated from the solution of this consistency condition achieve asymptotic social optimality.
Our first step is to evaluate precisely the impact of a single agent’s strategy on the expected social cost $J_{soc}$. To do this, we divide $J_{soc}$ into three parts, one of them depends only on agent $i$’s strategy, while the others are independent of it. Specifically, we state that:
Theorem 4.1.
If the strategies $\{\check{v}_{t}^{i}\}_{i=1}^{N}$ optimize $J_{soc}$, then for any $i=1,\ldots,N$, $\check{v}_{t}^{i}$ is the solution for minimizing control problem (4.1)-(4.2) when other agents’ maintain their control $\check{v}^{-i}$.
Proof.
For $i=1,\ldots,N$, let $J_{soc}=Y_{1}^{i}+Y_{2}^{i}$, where
$Y_{1}^{i}:=\frac{1}{N}\sum_{k\neq i}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{k}-\kappa)^{2}+\frac{c}{2}(v_{t}^{k})^{2}\right)dt+\frac{\gamma}{2}\left(X_{T}^{k}-\zeta\right)^{2}\right],$
$Y_{2}^{i}:=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+\Big(\sum_{k=1}^{N}v_{t}^{k}\Big)P_{t}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right].$
Note that the price process $P$ is only partially influenced by strategy $v^{i}$. Solving (2.2), we obtain
$P_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds,$
Using the decomposition $Q_{t}=\frac{1}{N}v_{t}^{i}+\frac{1}{N}\sum_{k\neq i}v_{t}^{k}$, we further decompose $Y_{2}^{i}=Y_{3}^{i}+Y_{4}^{i}$, where the components are given by
$Y_{3}^{i}$ $=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{k\neq i}v_{t}^{k}\right)\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\left(\sum_{k\neq i}v_{s}^{k}\right)ds\right)dt\right],$
$Y_{4}^{i}$ $=\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}+\frac{c}{2}{(v_{t}^{i})}^{2}+v_{t}^{i}P_{t}\right)dt+\frac{\gamma}{2}\left(X_{T}^{i}-\zeta\right)^{2}\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{k\neq i}v_{t}^{k}\right)\left(-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds\right)dt\right].$
We thus observe that $J_{soc}=Y_{1}^{i}+Y_{3}^{i}+Y_{4}^{i}$, and both $Y_{1}^{i}$ and $Y_{3}^{i}$ are independent of $v^{i}$. Since $Y_{4}^{i}=\frac{1}{N}\check{J}_{i}(v^{i})$, we conclude that if $(\check{v}_{t}^{1},...,\check{v}_{t}^{N})$ minimizes $J_{soc}$, then for $i=1,\ldots,N$, the strategy $\check{v}_{t}^{i}$ minimizes $\check{J}_{i}(v^{i})$. The proof is complete. ∎
Subsequently, we solve the optimal control problem (4.4). From (4.2), we obtain
$du_{t}^{i}=(-\alpha u_{t}^{i}+\alpha v_{t}^{i})dt,\quad u_{0}^{i}=0.$
Therefore, we consider the system
$\left\{\begin{aligned} &dX_{s}=v_{s}ds+\sigma_{i}dW_{s},\quad X_{t}=x,\\ &du_{s}=(-\alpha u_{s}+\alpha v_{s})ds,\quad u_{t}=u,\quad\forall t\leq s\leq T,\end{aligned}\right$. (4.5)
and we define the admissible strategy set for $v$ as
[…]
By dynamic programming principle, the value function
$V_{i}(t,x,u)=\inf_{v^{i}\in\mathcal{C}_{i}}\mathbb{E}\left[\int_{t}^{T}\Big(\frac{\eta}{2}(X_{s}^{i}-\kappa)^{2}+\frac{c}{2}(v_{s}^{i})^{2}-\bar{q}(s)u_{s}^{i}+v_{s}^{i}\bar{p}(s)\Big)ds+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
satisfies the following HJB equation
$\frac{\partial V_{i}}{\partial t}+\inf_{v^{i}\in\mathcal{C}_{i}}\left\{\frac{\partial V_{i}}{\partial x}v_{t}^{i}+\frac{\partial V_{i}}{\partial u}(\alpha v_{t}^{i}-\alpha u)+\frac{c}{2}(v_{t}^{i})^{2}+v_{t}^{i}\bar{p}(t)\right\}-\bar{q}(t)u+\frac{\sigma_{i}^{2}}{2}\frac{\partial^{2}V_{i}}{\partial^{2}x}+\frac{\eta}{2}(x-\kappa)^{2}=0.$
with terminal condition $V_{i}(T,x,u)=\frac{\gamma}{2}(x-\zeta)^{2}$. We adopt the ansatz
$V_{i}(t,x,u)=a_{i}(t)x^{2}+b_{i}(t)x+l_{i}(t)u+f_{i}(t),$
which reduces the HJB equation to
$a^{\prime}_{i}(t)x^{2}+b^{\prime}_{i}(t)x+f_{i}^{\prime}(t)+l^{\prime}_{i}(t)u+\inf_{v^{i}\in\mathcal{A}}\Big\{(2a_{i}(t)x+b_{i}(t))v_{t}^{i}+\alpha l_{i}(t)v_{t}^{i}-\alpha ul_{i}(t)$
$\quad+\frac{c}{2}{(v_{t}^{i})}^{2}+v_{t}^{i}\bar{p}(t)\Big\}-\bar{q}(t)u+\sigma_{i}^{2}a(t)+\frac{\eta}{2}(x-\kappa)^{2}=0.$
From this, we obtain the explicit optimal control
$\check{v}_{t}^{i}=-\frac{2a_{i}(t)\check{X}_{t}^{i}+b_{i}(t)+\bar{p}(t)+\alpha l_{i}(t)}{c}.$ (4.6)
Substituting (4.6) into the HJB equation yields the following ODE system:
$\qquad\qquad\qquad a^{\prime}_{i}(t)+\frac{\eta}{2}-\frac{2a^{2}_{i}(t)}{c}=0,$ $a_{i}(T)=\frac{\gamma}{2},$ (4.7)
$\qquad\qquad\qquad b^{\prime}_{i}(t)-\eta\kappa-\frac{2a_{i}(t)(\bar{p}(t)+b_{i}(t)+\alpha l_{i}(t))}{c}=0,$ $b_{i}(T)=-\gamma\zeta,$ (4.8)
$\qquad\qquad\qquad f^{\prime}_{i}(t)+\sigma_{i}^{2}a_{i}(t)+\frac{\eta\kappa^{2}}{2}-\frac{(b_{i}(t)+\bar{p}(t)+\alpha l_{i}(t))^{2}}{2c}=0,$ $f_{i}(T)=\frac{\gamma\zeta^{2}}{2},$ (4.9)
$\qquad\qquad\qquad l^{\prime}_{i}(t)-\alpha l_{i}(t)-\bar{q}(t)=0,$ $l_{i}(T)=0.$ (4.10)
We assert that (4.7)–(4.10) admit a unique solution on $[0,T]$. The proof is nearly identical to that of section 3, so we only provide a brief outline. First, (4.7) has a unique solution on $[0,T]$ by section 3. Next, as in the proof of section 3, both (4.10) and (4.8) admit unique solutions on $[0,T]$. Finally, uniqueness of the solution to (4.9) follows immediately by integrating the equation from $t$ to $T$. Since $a_{i}(t),\ b_{i}(t)$ and $l_{i}(t)$ are identical for all agents, we abbreviate them as a(t),b(t)a(t),\ b(t) and $l(t)$, respectively. Following the same argument as in Theorem 3.2, one can verify the optimality of (4.6), although such verification is not required for establishing the results in Theorem 4.5.
As in the analysis of the non-cooperative game, when all agents adopt the strategy in (4.6), the empirical mean of $X_{t}^{i}$, denoted by $\overline{X}_{t}$, can be approximated by a function $(\bar{x}(t))_{t\in[0,T]}$, whose dynamics satisfy
$d\bar{x}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}dt,\quad\bar{x}(0)=\bar{x}_{0}^{N}.$
This implies that the requirement in step 3 is equivalent to $\bar{q}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}$. Combining this with (4.3), (4.8), (4.10), we observe that the requirement is equivalent to ensuring that the system
$\frac{d}{dt}\begin{pmatrix}\bar{p}(t)\\ \bar{x}(t)\\ b(t)\\ l(t)\end{pmatrix}=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}(t)\\ \bar{x}(t)\\ b(t)\\ l(t)\end{pmatrix}+\begin{pmatrix}\alpha\beta\\ 0\\ \eta\kappa\\ 0\end{pmatrix}$, (4.11)
together with the associated initial and terminal conditions,
$\begin{pmatrix}\bar{p}(0),\ \bar{x}(0),\ b(T),\ l(T)\end{pmatrix}^{\top}=\begin{pmatrix}p_{0},\ \overline{x}_{0}^{N},\ -\gamma\zeta,\ 0\end{pmatrix}^{\top}$, (4.12)
admits a solution on $[0,T]$.
As before, the main challenge lies in the boundary conditions being imposed at different time points. To overcome this, we first construct two processes whose initial values are both specified at $t=0$, as described below.
Proposition 4.2.
Equation systems
$\frac{d}{dt}\begin{pmatrix}\bar{p}_{1}(t)\\ \bar{x}_{1}(t)\\ {b}_{1}(t)\\ {l}_{1}(t)\end{pmatrix}$ $=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}_{1}(t)\\ \bar{x}_{1}(t)\\ {b}_{1}(t)\\ {l}_{1}(t)\end{pmatrix}$, (4.13)
and
$\frac{d}{dt}\begin{pmatrix}\bar{p}_{2}(t)\\ \bar{x}_{2}(t)\\ {b}_{2}(t)\\ {l}_{2}(t)\end{pmatrix}$ $=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}\begin{pmatrix}\bar{p}_{2}(t)\\ \bar{x}_{2}(t)\\ {b}_{2}(t)\\ {l}_{2}(t)\end{pmatrix}$, (4.14)
associated with initial conditions $(\bar{p}_{1}(0),\ \bar{x}_{1}(0),\ {b}_{1}(0),\ {l}_{1}(0))^{\top}=(0,0,1,0)^{\top}$ and $(\bar{p}_{2}(0),\ \bar{x}_{2}(0),\ {b}_{2}(0),\ {l}_{2}(0))^{\top}=(0,0,0,1)^{\top}$, admit a unique solution, denoted by $(\phi_{1}^{*}(t))_{t\in[0,T]}=((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $(\phi_{2}^{*}(t))_{t\in[0,T]}$ $((\bar{p}_{2}(t),\ \bar{x}_{2}(t),\ b_{2}(t),\ l_{2}(t))^{\top})_{t\in[0,T]}$ respectively.
Next, we show that when the initial values of (4.11) are specified only at $t=0$, the system admits a unique solution. This implies that the solution to (4.11) is uniquely determined by $b(0)$ and $l(0)$, once given $\bar{p}(0)=p_{0}$ and $\bar{x}(0)=\bar{x}_{0}^{N}$.
Proposition 4.3.
For $\forall b_{0},\ l_{0}\in\mathbb{R}$, the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a unique solution that satisfies (4.11), denoted by
$(\phi_{b_{0},l_{0}}^{*}(t))_{t\in[0,T]}=((\bar{p}_{b_{0},l_{0}}(t),\ \bar{x}_{b_{0},l_{0}}(t),\ b_{b_{0},l_{0}}(t),\ l_{b_{0},l_{0}}(t))^{\top})_{t\in[0,T]}.$
We now introduce a sufficient condition ensuring the existence and uniqueness of solutions to (4.11)–(4.12), stated in Assumption $\boldsymbol{(A_{3})}$.
Assumption $\boldsymbol{(A_{3}):}$ The systems $((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $((\bar{p}_{2}(t),\ \bar{x}_{2}(t)$, $b_{2}(t),$ $l_{2}(t))^{\top})_{t\in[0,T]}$, which are given by (4.13) and (4.14) respectively, satisfy that $({b}_{1}(T),\ {l}_{1}(T))^{\top}$, $({b}_{2}(T),\ {l}_{2}(T))^{\top}$ are linearly independent vectors.
section 4 below illustrates that Assumption $\boldsymbol{(A_{3})}$ is sufficient for ensuring the existence of solution to (4.11) and (4.12).
Proposition 4.4.
If Assumption $\boldsymbol{(A_{3})}$ holds, then the existence and uniqueness of solution to (4.11) and (4.12) also hold.
As in other works on mean field games, a coercivity condition is required to prevent the absence of an optimal or $\epsilon$-optimal control. However, this condition need not hold for arbitrary positive parameters $c$ and $\eta$. To ensure coercivity, we impose the following assumption.
Assumption $\boldsymbol{(A_{4}):}$ The parameters $c$ and $\eta$ satisfy that $(c-2)\eta>\alpha^{2}$.
Remark 3.
To illustrate the necessity of this assumption, consider the case in which $v_{t}^{i}$ remains being a large constant. Then, the term $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}(v_{t}^{i}P_{t})dt\right]$ can become a large negative number, potentially dominating the positive contributions from $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\left(\frac{c}{2}{(v_{t}^{i})}^{2}+\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}\right)dt\right]$ when $c$ and $\eta$ are sufficiently small. Consequently, $J_{soc}$ may approach -$\infty$ as the control norm increases, violating coercivity. Since the coefficient in front of $v_{t}^{i}P_{t}$ in $J_{soc}$ is 1, requiring $\frac{c}{2}>1$ and a proper lower bound for $\eta$ is sufficient to guarantee coercivity, as stated in Assumption $\boldsymbol{(A_{4})}$.
Under these assumptions, we construct the process described in Theorem 4.5 and show that it yields a set of $\epsilon$-optimal strategies.
Theorem 4.5.
Assume that Assumption $\boldsymbol{(A_{2})}$, $\boldsymbol{(A_{3})}$ and $\boldsymbol{(A_{4})}$ hold, and the unique solution for (4.11) and (4.12) is $(\bar{p}(t),\ \bar{x}(t),\ b(t),\ l(t))^{\top}$. We construct the system
$\left\{\begin{aligned} &d\check{X}_{t}^{i}=\check{v}_{t}^{i}dt+\sigma_{i}dW_{t}^{i},\quad\check{X}_{0}^{i}=x_{0}^{i},\\ &\check{v}_{t}^{i}=-\frac{2a(t)\check{X}_{t}^{i}+{b}(t)+\bar{p}(t)+\alpha{l}(t)}{c},\\ &\check{Q}_{t}=\frac{1}{N}\sum_{i=1}^{N}\check{v}_{t}^{i},\\ &d\check{P}_{t}=\alpha(\beta-\check{P}_{t}-\check{Q}_{t})dt,\quad\check{P}_{0}=p_{0}.\end{aligned}\right$. (4.15)
Then, the set of strategies $\{\check{v}^{i}\}_{1\leq i\leq N}$ is a set of $\epsilon$-optimal strategies for $J_{soc}$. Moreover, specific calculation shows that under $\{\check{v}_{t}^{i}\}_{i=1}^{N}$, $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$ in (2.5).
Suppose there exists another admissible control profile $\boldsymbol{v}=\{v^{i}\}_{1\leq i\leq N}$, and let the corresponding processes evolve according to
$\left\{\begin{aligned} &d{X_{t}^{i}}={v_{t}^{i}}dt+\sigma_{i}dW_{t}^{i},\quad X_{t}^{i}=x_{0}^{i},\\ &Q_{t}=\frac{1}{N}\sum_{i=1}^{N}{v_{t}^{i}},\\ &dP_{t}=\alpha(\beta-P_{t}-Q_{t})dt,\quad P_{0}=p_{0}.\end{aligned}\right$.
Our objective is to show that (2.5) holds, or equivalently
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}})\geq-\epsilon_{N},$ (4.16)
where $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$. To establish this, we introduce two lemmas. Lemma 4.2 characterizes the coercivity condition. To facilitate a clearer proof, we first present Lemma 4.1.
Lemma 4.1.
For any $i=1,2,\ldots,N$ and $v\in\mathcal{A}_{i}$, we construct the process
$x_{t}=\begin{pmatrix}\tilde{X}_{t}\\ \tilde{P}_{t}\end{pmatrix},\quad\frac{d}{dt}x_{t}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}x_{t}+\begin{pmatrix}1\\ -\alpha\end{pmatrix}v_{t},\quad x_{0}=\begin{pmatrix}0\\ 0\end{pmatrix}$.
Then, the following result
$\mathbb{E}\left[\int_{0}^{T}v_{t}\Big(v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}\Big)dt\right]\geq 0$
holds.
We now proceed to formulate the coercivity condition rigorously, as stated in Lemma 4.2.
Lemma 4.2.
Assume that Assumptions $\boldsymbol{(A_{2})}$ and $\boldsymbol{(A_{4})}$ hold. Then there exists constants $\epsilon>0,\ \epsilon^{\prime}>0$, and $C\in\mathbb{R}$, all independent of $N$ and $i$, such that
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}({\epsilon}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2})dt\right]+C.$
Having established the necessary preparations, we now prove Theorem 4.5.
Proof of Theorem 4.5.
Let
$\Delta X_{t}^{i}=X_{t}^{i}-\check{X}_{t}^{i},\quad\Delta v_{t}^{i}=v_{t}^{i}-\check{v}_{t}^{i},\quad\Delta P_{t}=P_{t}-\check{P}_{t},\quad\Delta Q_{t}=Q_{t}-\check{Q}_{t},\quad\Delta u_{t}^{i}=u_{t}^{i}-\check{u}_{t}^{i},$
where $u^{i}$ and $\check{u}^{i}$ are given by
$du_{t}^{i}=-\alpha u_{t}^{i}dt+\alpha v_{t}^{i}dt,\quad u_{0}^{i}=0,$
$d\check{u}_{t}^{i}=-\alpha\check{u}_{t}^{i}dt+\alpha\check{v}_{t}^{i}dt,\quad\check{u}_{0}^{i}=0.$
Let $(\check{X}_{t}^{i},\ \check{P}_{t},\ \check{Q}_{t})$ denote the processes induced by the optimal control in (4.15). Under this notation, the difference in the social cost can be written as
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}})$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(\check{X}_{t}^{i}+\Delta X_{t}^{i}-\kappa\right)^{2}-\frac{\eta}{2}\left(\check{X}_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}\left(\check{v}_{t}^{i}+\Delta v_{t}^{i}\right)^{2}-\frac{c}{2}({\check{v}_{t}^{i})}^{2}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\check{v}_{t}^{i}+\Delta{v_{t}^{i}})(\check{P}_{t}+\Delta{P}_{t})-\check{v}_{t}^{i}\check{P}_{t}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\frac{\gamma}{2}\left(\check{X}_{T}^{i}+\Delta X_{T}^{i}-\zeta\right)^{2}-\frac{\gamma}{2}\left(\check{X}_{T}^{i}-\zeta\right)^{2}\right].$ (4.17)
Applying Itô’s formula, we convert the terminal contribution into the integral form,
$\mathbb{E}\bigg[\frac{\gamma}{2}\left(\check{X}_{T}^{i}+\Delta X_{T}^{i}-\zeta\right)^{2}-\frac{\gamma}{2}\left(\check{X}_{T}^{i}-\zeta\right)^{2}\bigg]$
$=\mathbb{E}\left[\int_{0}^{T}\left(a^{\prime}(t)(\Delta X_{t}^{i})^{2}+2a^{\prime}(t)\Delta X_{t}^{i}\check{X}_{t}^{i}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(2a(t)\Delta X_{t}^{i}\check{v}_{t}^{i}+2a(t)\Delta v_{t}^{i}\check{X}_{t}^{i}+{b}^{\prime}(t)\Delta X_{t}^{i}+{b}(t)\Delta v_{t}^{i}\right)dt\right].$ (4.18)
Substituting this into (4.17) gives
$J_{soc}(\boldsymbol{v})-J_{soc}(\check{\boldsymbol{v}}):=I_{1}+I_{2},$
where $I_{1}$ and $I_{2}$ are defined by
$I_{1}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\bigg[\int_{0}^{T}\left((\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)+\Delta{X_{t}^{i}}\check{X}_{t}^{i}\left(\eta+2a^{\prime}(t)-\frac{(2a(t))^{2}}{c}\right)\right)dt\bigg]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\Delta{X_{t}^{i}}\left(-\eta\kappa+{b}^{\prime}(t)-\frac{2a(t)(\bar{p}(t)+\alpha{l}(t)+{b}(t))}{c}\right)+\Delta v_{t}^{i}\left(-\alpha l(t)\right)\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}+\bar{q}(t)\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right]$
$=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)-\Delta v_{t}^{i}\alpha{l}(t)+\frac{c}{2}(\Delta v_{t}^{i})^{2}\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}+\Delta v_{t}^{i}\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right],$ (4.19)
and
$I_{2}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left((\check{v}_{t}^{i}-\bar{q}(t))\Delta P_{t}+\Delta Q_{t}(\check{P}_{t}-\bar{p}(t)\right)dt\right].$
We now provide a more detailed explanation of how we derive $I_{1}$ and what $I_{2}$ characterizes. To derive $I_{1}$, we approximate $\sum_{i=1}^{N}\check{v}_{t}^{i}\Delta P_{t}$ by $\sum_{i=1}^{N}\overline{q}(t)\Delta P_{t}$ and approximate $\sum_{i=1}^{N}\check{P}_{t}\Delta v_{t}^{i}$ by $\sum_{i=1}^{N}\overline{p}(t)\Delta v_{t}^{i}$. We then use (4.18) to rewrite the terminal terms, and apply the optimal control formula (4.15) to eliminate $\check{v}_{t}^{i}$. This yields $I_{1}$, which represents the leading-order change in $J_{soc}$ caused by deviating from the proposed strategies. The second part, $I_{2}$, captures the approximation error due to replacing $(\check{v}_{t}^{i},\check{P}_{t})$ with $(\bar{q},\bar{p})$.
Following exactly the same argument as in the proof of Lemma 3.1, we obtain the existence of an $O(1)$ constant $K^{\prime\prime}$ such that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\check{X}_{t}^{i}|^{2}\right]\vee\mathbb{E}\left[\sup_{0\leq t\leq T}|\check{Q}_{t}|^{2}\right]\leq K^{\prime\prime},\ \mathbb{E}\left[\int_{0}^{T}({\check{v}_{t}^{i}})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\check{v}_{t}^{i})^{2}dt\right]\leq K^{\prime\prime},\ i=1,\ldots,N,$
and
$|J_{i}(\check{v}^{i},\check{\boldsymbol{v}}^{-i})|\leq K^{\prime\prime},\quad i=1,2,\ldots,N,$
which further implies
$J_{soc}(\check{\boldsymbol{v}})\leq K^{\prime\prime}.$
By the coercivity condition in Lemma 4.2, there exists $O(1)$ constant $M_{1}>0$ such that, if $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]>M_{1}$, we have
$J_{soc}(\boldsymbol{v})\geq J_{soc}(\check{\boldsymbol{v}}).$
This inequality directly gives us the desired result. Thus, it suffices to establish the desired inequality (4.16) under the regime $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]\leq M_{1}$. Indeed, we will show that
$I_{1}\geq 0,\quad I_{2}=O\bigg(\frac{1}{\sqrt{N}}\bigg).$
We first prove the claim for the term $I_{1}$. The next step relies on the identity
$\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=0.$ (4.20)
We sketch the argument. Applying Itô’s formula and using the evolution of l{l} in (4.10), we obtain
$0$ $=\mathbb{E}\left[{l}(T)\Delta u_{T}^{i}\right]=\mathbb{E}\left[\int_{0}^{T}\left({l}^{\prime}(t)\Delta u_{t}^{i}+{l}(t)(-\alpha\Delta u_{t}^{i}+\alpha\Delta v_{t}^{i})\right)dt\right]$
$=\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta u_{t}^{i}dt+\alpha{l}(t)\Delta v_{t}^{i}\right)dt\right],$ (4.21)
with $\Delta{P_{t}}$ and $\Delta u_{t}^{i}$ satisfying,
$d\Delta{P_{t}}=\alpha(-\Delta{P_{t}}-\Delta Q_{t})dt,\quad\Delta{P_{0}}=0,$ (4.22)
$d\Delta u_{t}^{i}=\alpha(-\Delta u_{t}^{i}+\Delta v_{t}^{i})dt,\quad\Delta u_{0}^{i}=0.$ (4.23)
Since $\Delta Q_{t}$ is the average of $\Delta v_{t}^{i}$, it follows from these dynamics that
$\frac{1}{N}\sum_{i=1}^{N}\Delta u_{t}^{i}=-\Delta P_{t},$ (4.24)
and then
$\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\bar{q}(t)\Delta P_{t}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(-\bar{q}(t)\Delta u_{t}^{i}-\Delta v_{t}^{i}\alpha{l}(t)\right)dt\right]=0,$
where the first and the second step are derived from (4.24) and (4) respectively. This leads directly to (4.20). Substituting (4.20) into the expression for $I_{1}$ and applying Itô’s formula again, we obtain
$I_{1}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\sum_{i=1}^{N}(\Delta{X_{t}^{i}})^{2}\left(\frac{\eta}{2}+a^{\prime}(t)\right)+\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}+2a(t)\Delta v_{t}^{i}\Delta X_{t}^{i}\right)dt\right].$
Applying Itô’s formula once more, we obtain
$\mathbb{E}\left[a(T)(\Delta X_{T}^{i})^{2}\right]=\mathbb{E}\left[\int_{0}^{T}\left(a^{\prime}(t)(\Delta X_{t}^{i})^{2}+2a(t)\Delta X_{t}^{i}\Delta v_{t}^{i}\right)dt\right].$
This further simplifies the expression of $I_{1}$ into
$I_{1}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(\Delta{X_{t}^{i}})^{2}+\frac{c}{2}(\Delta v_{t}^{i})^{2}+\Delta v_{t}^{i}\Delta P_{t}\right)dt+\frac{\gamma}{2}(\Delta X_{T}^{i})^{2}\right].$
Let
$\overline{\Delta X_{t}}:=\frac{1}{N}\sum_{i=1}^{N}\Delta X_{t}^{i}$
be the averaged deviation process. Note that
$\left\{\begin{aligned} &d\overline{\Delta X_{t}}=\Delta Q_{t}dt,\\ &d\Delta P_{t}=-\alpha\Delta P_{t}dt-\alpha\Delta Q_{t}dt,\\ &\overline{\Delta X_{0}}=0,\quad\Delta P_{0}=0.\end{aligned}\right$.
By Lemma 4.1, we conclude that
$\mathbb{E}\left[\int_{0}^{T}\left(\Delta Q_{t}+\Delta P_{t}+\alpha\overline{\Delta X_{t}}\right){\Delta Q_{t}}dt\right]\geq 0.$
In order to use this inequality, we invoke Assumption $\boldsymbol{(A_{4})}$, and apply fundamental inequality to obtain
$\frac{\eta}{2}\left(\overline{\Delta X_{t}}\right)^{2}-\alpha\overline{\Delta X_{t}}\Delta Q_{t}+\frac{c-2}{2}\Delta Q_{t}^{2}\geq 0.$
Subsequently, the convexity of $\frac{\eta}{2}x^{2}$ leads to
$I_{1}$ $\geq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(\overline{\Delta{X_{t}}}\right)^{2}+\frac{c}{2}\left(\Delta Q_{t}\right)^{2}+\Delta v_{t}^{i}\Delta P_{t}\right)dt\right]$
$=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\Delta Q_{t}(\Delta P_{t}+\alpha\overline{\Delta X_{t}}+\Delta Q_{t})\right)dt\right]$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}(\overline{\Delta X_{t}})^{2}-\alpha\overline{\Delta X_{t}}\Delta Q_{t}+\frac{c-2}{2}(\Delta Q_{t})^{2}\right)dt\right],$ (4.25)
where Jensen’s inequality is used to replace $\sum_{i=1}^{N}(\Delta X_{t}^{i})^{2}$ by $N\overline{\Delta X_{t}}^{2}$, and $\sum_{i=1}^{N}(\Delta v_{t}^{i})^{2}$ by $N\Delta Q_{t}^{2}$. Since the right hand side of (4.25) is nonnegative, we conclude
$I_{1}\geq 0.$
The proof for $I_{1}$ is completed, and it is straightforward to prove the estimate for $I_{2}$. Similar as in Theorem 3.6’s proof, we have
$\mathbb{E}\left[\int_{0}^{T}\bigg(\frac{1}{N}\sum_{i=1}^{N}\big(\check{v}_{t}^{i}-\bar{q}(t)\big)\bigg)^{2}dt\right]=O\Big(\frac{1}{N}\Big).$
Using the Cauchy–Schwarz and triangle inequalities, together with the bound $\|Q\|^{2}\leq M_{1}$, we obtain
$\|\Delta Q\|^{2}\leq O(1)(\|Q\|+\|\check{Q}\|)^{2}\leq O(1).$
Moreover, solving (4.22) and applying Cauchy–Schwarz again yields
$\Delta P_{t}=-\alpha\int_{0}^{t}e^{-\alpha(t-s)}\Delta Q_{s}ds,$
and
$|\Delta P_{t}|^{2}\leq\alpha^{2}\|\Delta Q\|^{2}\left(\int_{0}^{t}e^{-2\alpha(t-s)}ds\right)\leq O(1)\|\Delta Q\|^{2},\quad\|\Delta P\|^{2}\leq O(1)\|\Delta Q\|^{2}.$
Thus,
$\frac{1}{N}\left|\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\big(\check{v}_{t}^{i}-\bar{q}(t)\big)\Delta P_{t}dt\right]\right|\leq\frac{1}{N}\|\Delta P\|\cdot\left\|\sum_{i=1}^{N}(\check{v}^{i}-\bar{q})\right\|\leq O\bigg(\frac{1}{\sqrt{N}}\bigg).$ (4.26)
Proceeding in the same way as in Theorem 3.6, we get
$\|\check{P}-\bar{p}\|=O\bigg(\frac{1}{\sqrt{N}}\bigg).$
Hence,
$\bigg|\mathbb{E}\left[\int_{0}^{T}\Delta Q_{t}(\check{P}_{t}-\bar{p}(t))dt\right]\bigg|\leq\|\Delta Q\|\cdot\|\check{P}-\bar{p}\|\leq O\bigg(\frac{1}{\sqrt{N}}\bigg).$ (4.27)
Therefore, combining (4.26) and (4.27), we conclude that
$I_{2}=O\bigg(\frac{1}{\sqrt{N}}\bigg).$
The proof is complete.
5 Numerical experiments
[…]
Then, we compare the discrepancy between the approximation functions and the true values under the non-cooperate case and the cooperate case through numerical experiments. As shown below, 1(a) and 1(b) compare the evolution of $\bar{p}$ and $P$, $\bar{q}$ and $Q$ in the non-cooperative game setting, respectively. 2(a) and 2(b) present the same comparisons under the cooperative game setting. From the plot, we can see that in the non-cooperate case, both curves in 1(a) exhibit a decreasing trend, and both curves in 1(b) display a $U$-shaped trajectory. However, the cooperative strategies not only lead to a steeper initial decline in the price, but also cause $Q$ to decrease monotonically. Besides, the cooperative case yields more stable and smoother trajectories, especially for the $Q$-related dynamics. This suggests that cooperation among agents can lead to more efficient and consistent outcomes in price and the average trading rate. Although the shapes of the $P$ and $Q$ curves differ between the figures in both cases, these figures all demonstrate that the mean-field approximation performs very well in both cooperative and non-cooperative settings. This property could be valuable in practical applications like smart grids or large-scale decentralized systems.
(a) Curves of $\bar{P}$ and $P$ in the non-cooperate case
(b) Curves of $\bar{Q}$ and $Q$ in the non-cooperate case
Figure 1: Non-cooperate case: evolution of $\bar{P},P$ and $\bar{Q},Q$
(a) Curves of $\bar{p}$ and $P$ in the cooperate case
(b) Curves of $\bar{q}$ and $Q$ in the cooperate case
Figure 2: Cooperate case: evolution of $\bar{p},P$ and $\bar{q},Q$
As shown in Figure 3, we also compare the evolution of the control variable $v$ for the first 10 agents under both cases. In 3(a), which corresponds to the non-cooperative case, the trajectories of $v^{i}$ appear more tightly clustered, with relatively smaller deviations from one another over time. In contrast, 3(b), which shows the cooperative case, exhibits greater divergence in the trajectories of $v^{i}$. Overall, the figures illustrate how cooperation introduces more variability in individual control paths, while non-cooperation leads to more uniform behavior across agents. This reflects the fact that, in a cooperative setting, agents can adjust their behaviors more flexibly to improve overall system performance.
(a) Evolution of $v^{i}$ for agent 1 to 10 (the non-cooperate case)
(b) Evolution of $v^{i}$ for agent 1 to 10 (the cooperate case)
Figure 3: Comparison of the evolution process of $v^{i}$ under both cases
6 Conclusion
This paper has developed a strategic framework for smart grids with many agents using mean field game methodology. Under sticky prices and finite time horizons, we analyze both the cooperative and non-cooperative settings. In the cooperative case, we derive strategies that minimize the expected social cost, while in the non-cooperative case, we construct approximate Nash equilibrium. Numerical experiments demonstrate the effectiveness of these approximations, offering insights for managing large-scale smart grid systems.
Appendix A Proofs of Auxiliary Lemmas
In this appendix, we collect the proofs of the auxiliary lemmas. We start with the proof of Lemma 3.1.
Proof of Lemma 3.1.
It follows from (3.12) that
$d\hat{X}_{t}^{i}=-\frac{\bar{P}(t)+2a(t)\hat{X}_{t}^{i}+B(t)}{c}dt+\sigma_{i}dW_{t}^{i},\quad\hat{X}_{0}^{i}=x_{0}^{i},$
which implies that
$d\overline{\hat{X}}_{t}=-\frac{\bar{P}(t)+2a(t)\overline{\hat{X}}_{t}+B(t)}{c}dt+\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}dW_{t}^{i},\quad\overline{\hat{X}}_{0}=\bar{x}_{0}^{N}.$
Gronwall Lemma, along with Assumption $\boldsymbol{(A_{2})}$, shows that there exists an $O(1)$ number $K$ such that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\hat{X}_{t}^{i}|^{2}\right]\leq K,\ \quad\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\overline{\hat{X}}_{t}\right|^{2}\right]\leq K,\ \ \forall i=1,\ldots,N.$
Therefore, from (3.12), there exists an $O(1)$ number (still denoted by $K$) that satisfies
$\mathbb{E}\left[\int_{0}^{T}(\hat{v}_{t}^{i})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}(\hat{v}_{t}^{i})^{2}dt\right]\leq K,\ \forall i=1,\ldots,N,\$
$\mathbb{E}\left[\int_{0}^{T}\hat{Q}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}\hat{Q}_{t}^{2}\right]\leq K,\ \forall i=1,\ldots,N.$ (A.1)
For any $i$, we may treat all other agents as a large population of which consists of N1N-1 agents. Applying the calculations from the previous calculations, we obtain
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N-1}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K.$
This leads to
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\frac{1}{N}\sum_{k\neq i}\hat{v}_{t}^{k}\right|^{2}\right]\leq K.$
To proceed, we require an exact expression of $\hat{P}$. Solving directly from (2.2) yields
$\hat{P}_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha\hat{Q}_{s}ds.$
It follows that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]$ $\leq 2\mathbb{E}\left[\sup_{0\leq t\leq T}\left\{\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds\right)^{2}\right\}\right]$
$\quad+2\mathbb{E}\left[\sup_{0\leq t\leq T}\left\{\left(\int_{0}^{t}e^{-\alpha(t-s)}\alpha\hat{Q}_{s}ds\right)^{2}\right\}\right]$
$\leq O(1)+2(\alpha T)^{2}\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{Q}_{t}}|^{2}\right]$
$\leq O(1),$
where we have used the bounds in (A). This shows that
$\mathbb{E}\left[\int_{0}^{T}(\hat{P}_{t})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq O(1).$
Moreover, (A) implies that
$\left|\mathbb{E}\left[\int_{0}^{T}\bigg(\frac{\eta}{2}(\hat{X}_{t}^{i}-\kappa)^{2}+\frac{c}{2}(\hat{v}_{t}^{i})^{2}\bigg)dt+\frac{\gamma}{2}(\hat{X}_{T}-\zeta)^{2}\right]\right|\leq O(1).$
The $O(1)$ bound for $\|\hat{v}\|$ and $\|\hat{P}\|$, together with Cauchy-Schwartz inequality, implies
$\left|\mathbb{E}\left[\int_{0}^{T}\hat{v}_{t}^{i}\hat{P}_{t}dt\right]\right|\leq\|\hat{v}^{i}\|\cdot\|\hat{P}\|\leq O(1).$
This establishes that
$|J_{i}(\hat{v}^{i},\hat{\boldsymbol{v}}^{-i})|\leq O(1),\quad i=1,\ldots,N.$
Let $K^{\prime}$ be an $O(1)$ number that exceeds $K$ and all the $O(1)$ numbers above. With this, the proof of Lemma 3.1 is complete. ∎
We now turn to the proofs of Lemma 3.2 and Lemma 3.3.
Proof of Lemma 3.2.
Note that $v\in\mathcal{A}_{i}$ implies
$\int_{0}^{T}v_{t}^{2}dt<\infty,\ a.s.$
Consequently, we restrict attention to events for which the above integral is finite. Under such events, the integral in Lemma 3.2 exists and is finite. To present the proof clearly, we introduce the following notation:
$A_{1}=\begin{pmatrix}0&0\\ 0&{-\alpha}\end{pmatrix}\preceq O,\quad B_{1}=\begin{pmatrix}1\\ -\frac{\alpha}{N}\end{pmatrix},\quad C_{1}=\begin{pmatrix}{\epsilon}_{1}^{*}&1\end{pmatrix},\quad D_{1}={\epsilon}_{2}^{*},$
$x_{t}=\begin{pmatrix}X_{t}^{*}\\ P_{t}^{*}\end{pmatrix},\quad a^{*}={\epsilon}_{1}^{*}>0,\quad b^{*}=\frac{1}{2\alpha D_{1}}>0,\quad P_{1}=P_{1}^{\top}=diag{(a^{*},b^{*})}\succ O.$
Recalling (3.14), we define
$y_{t}:=\epsilon_{1}^{*}X_{t}^{*}+P_{t}^{*}+{\epsilon}_{2}^{*}v_{t}=C_{1}x_{t}+D_{1}v_{t}.$
Our objective is to apply Positive Real Lemma (Section 2.7.2, Boyd et al. (1994)) to the system. To proceed, we denote the LMI for the system by
$L=\begin{pmatrix}A_{1}^{\top}P_{1}+P_{1}A_{1}&P_{1}B_{1}-C_{1}^{\top}\\ B_{1}^{\top}P_{1}-C_{1}&-D_{1}-D_{1}^{\top}\end{pmatrix}=\begin{pmatrix}0&0&a^{*}-\epsilon_{1}^{*}\\ 0&-2{\alpha b^{*}}&-\frac{b^{*}\alpha}{N}-1\\ a^{*}-\epsilon_{1}^{*}&-\frac{b^{*}\alpha}{N}-1&-2D_{1}\end{pmatrix}$.
We investigate the conditions under which the LMI is semi-negative definite. By plugging in the exact value of $a^{*}$ and $b^{*}$, we discover that
$L\preceq O\iff O\preceq\begin{pmatrix}0&0&-a^{*}+\epsilon_{1}^{*}\\ 0&2{\alpha b^{*}}&\frac{b^{*}\alpha}{N}+1\\ -a^{*}+\epsilon_{1}^{*}&\frac{b^{*}\alpha}{N}+1&2D_{1}\end{pmatrix}=\begin{pmatrix}0&0&0\\ 0&\frac{1}{\epsilon_{2}^{*}}&\frac{1}{2\epsilon_{2}^{*}N}+1\\ 0&\frac{1}{2\epsilon_{2}^{*}N}+1&2\epsilon_{2}^{*}\end{pmatrix}$.
Therefore,
$L\preceq O\iff\text{det}\begin{pmatrix}\frac{1}{\epsilon_{2}^{*}}&\frac{1}{2\epsilon_{2}^{*}N}+1\\ \frac{1}{2\epsilon_{2}^{*}N}+1&2\epsilon_{2}^{*}\end{pmatrix}\geq 0\iff N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}.$
Since we assume $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, the LMI is semi-negative definite. Positive Real Lemma implies
$\int_{0}^{T}v_{t}(C_{1}x_{t}+D_{1}v_{t})dt\geq 0.$
Substituting the explicit forms of $C_{1}$ and $D_{1}$ and taking the expectation, we immediately obtain that Lemma 3.2 holds.
Proof of Lemma 3.3.
Fix positive numbers $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$ such that
$\epsilon_{2}=\frac{c}{8},\quad\epsilon_{3}<\frac{\eta}{4},\quad\epsilon_{4}+\epsilon_{5}<\frac{c}{8},\quad\epsilon_{3}\epsilon_{4}=\epsilon_{1}^{2}.$
Our first step uses (2.2), which gives
$P_{t}=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\Big(\sum_{k\neq i}\hat{v}_{s}^{k}\Big)ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds.$
To evaluate how $P$ changes in response to $v$, we decompose it into two components, one dependent on $v$ and the other independent of $v$. As shown below, we set
$P^{\prime}_{t}:=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\left(\sum_{k\neq i}\hat{v}_{s}^{k}\right)ds,$
$P_{t}^{\prime\prime}:=-\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}v_{s}^{i}ds.$
We naturally have $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$. The above expression for $P^{\prime\prime}_{t}$ is equivalently characterized by the dynamics $dP_{t}^{\prime\prime}=-\alpha P_{t}^{\prime\prime}-\frac{\alpha}{N}v_{t}^{i}$, $P^{\prime\prime}_{0}=0$. Both processes $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ are driven by $v^{i}$, and their coefficients match those in Lemma 3.2. However, applying the lemma directly requires that both initial conditions be zero, which does not hold in general. Thus, we take a detour and construct the process
$d\dot{X}_{t}^{i}=v_{t}^{i}dt,\ \dot{X}_{0}^{i}=0.$
Note that the dynamics and initial value of $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ coincide with those in Lemma 3.2. Therefore, we substitute $({X}_{t}^{*},\ P_{t}^{*})^{\top}$ with $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ in Lemma 3.2, and obtain that if $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$, then
$\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}(\epsilon_{1}\dot{X}_{t}^{i}+P^{\prime\prime}_{t}+\epsilon_{2}v_{t}^{i})dt\right]\geq 0.$ (A.2)
Combining $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$ and (A.2), we have, for all $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ $=\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}v_{t}^{i}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right]$
$\geq\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}^{\prime\prime}v_{t}^{i}-|v_{t}^{i}P_{t}^{\prime}|\right)dt\right]$
$=\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}\right)dt\right]+\mathbb{E}\left[\int_{0}^{T}v_{t}^{i}\left(\epsilon_{1}\dot{X}_{t}^{i}+P^{\prime\prime}_{t}+\epsilon_{2}v_{t}^{i}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{5}(v_{t}^{i})^{2}-|P_{t}^{\prime}v_{t}^{i}|\right)dt\right]$
$\geq\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}\right)dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{5}(v_{t}^{i})^{2}-|P_{t}^{\prime}v_{t}^{i}|\right)dt\right].$ (A.3)
From Lemma 3.1, direct calculations and the Cauchy–Schwarz inequality yield
$\mathbb{E}\left[\sup_{t\in[0,T]}(P^{\prime}_{t})^{2}\right]$ $\leq 2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds\right)^{2}\right\}\right]$
$\quad+2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\bigg(\sum_{k\neq i}\hat{v}_{s}^{k}\bigg)ds\right)^{2}\right\}\right]$
$\leq O(1),$
and hence
$\mathbb{E}\left[\int_{0}^{T}(P^{\prime}_{t})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{t\in[0,T]}(P^{\prime}_{t})^{2}\right]\leq O(1).$
The fundamental inequality gives $|P^{\prime}_{t}v_{t}^{i}|\leq\frac{\epsilon_{5}}{2}(v_{t}^{i})^{2}+\frac{1}{2\epsilon_{5}}(P^{\prime}_{t})^{2}$, leading to
$\mathbb{E}\left[\int_{0}^{T}\big(\epsilon_{5}(v_{t}^{i})^{2}-|P^{\prime}_{t}v_{t}^{i}|\big)dt\right]\geq\mathbb{E}\left[\int_{0}^{T}(\epsilon_{5}(v_{t}^{i})^{2}-\frac{1}{2}\epsilon_{5}(v_{t}^{i})^{2}-O(1)(P^{\prime}_{t})^{2})dt\right]\geq O(1).$
This simplifies (A) into the following inequality, which holds once $N$ is larger than $\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})\geq\mathbb{E}\left[\int_{0}^{T}\bigg(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}\bigg)dt\right]+O(1).$
We now control the possible negative cross-term $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}$ using suitable multiples of $(\dot{X}_{t}^{i})^{2}$ and $(v_{t}^{i})^{2}$. Since the inequality contains $X_{t}^{i}$, we rewrite $\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]$ in terms of $\dot{X}_{t}^{i}$. From $(X_{t}^{i}-\kappa)-(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)=\dot{X}_{t}^{i}$, we get
$\mathbb{E}[(\dot{X}_{t}^{i})^{2}]$ $\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]+2\mathbb{E}[(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)^{2}]$
$\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)]^{2}+4(x_{0}^{i}-\kappa)^{2}+4\sigma_{i}^{2}T,$
and hence
$\mathbb{E}\left[(X_{t}^{i}-\kappa)^{2}\right]\geq$ $\frac{1}{2}\mathbb{E}\left[(\dot{X}_{t}^{i})^{2}\right]+O(1),$
and, for $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}=\frac{4(\sqrt{2}+1)}{c}$
$J_{i}(v^{i},\hat{\boldsymbol{v}}^{-i})$ $\geq\mathbb{E}\left[\int_{0}^{T}\bigg(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{5}\right)(v_{t}^{i})^{2}-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}\bigg)dt\right]+O(1)$
$=\mathbb{E}\Bigg[\int_{0}^{T}\bigg(\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg)dt$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\epsilon_{3}(X_{t}^{i}-\kappa)^{2}\right)dt\right]+O(1)$
$\geq\mathbb{E}\Bigg[\int_{0}^{T}\bigg(\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg)dt$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\right)dt\right]+O(1),$
$\geq\mathbb{E}\Bigg[\int_{0}^{T}\bigg(\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg)dt+O(1),\$
where the last step relies on the restriction $\epsilon_{1}^{2}=\epsilon_{3}\epsilon_{4}$. Furthermore, applying the restrictions on $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$ yield the desired result. Therefore, the proof is complete.
The proofs of Lemma 4.1 and Lemma 4.2 are presented below.
Proof of Lemma 4.1.
For $v\in\mathcal{A}_{i}$, we impose the condition $\int_{0}^{T}v_{t}^{2}dt<\infty$, as in the proof of Lemma 3.2. We introduce the notion
$A_{2}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}\preceq O,\quad B_{2}=\begin{pmatrix}1\\ -\alpha\end{pmatrix},\quad C_{2}=\begin{pmatrix}\alpha&1\end{pmatrix},\quad D_{2}=1,$
and the output function
$\tilde{y}_{t}:=v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}=C_{2}\tilde{x}_{t}+D_{2},$
which allows us to rewrite the system as
$\frac{d}{dt}\tilde{x}_{t}=A_{2}\tilde{x}_{t}+B_{2},\quad\tilde{x}_{0}=(0,\ 0)^{\top}.$
We further introduce
$\tilde{a}=\alpha,\quad\tilde{b}=\frac{1}{\alpha},\quad P_{2}=diag{(\tilde{a},\tilde{b})}\succ O,$
and define the corresponding LMI by
$L_{2}=\begin{pmatrix}A_{2}^{\top}P_{2}+P_{2}A_{2}&P_{2}B_{2}-C_{2}^{\top}\\ B_{2}^{\top}P_{2}-C_{2}&-D_{2}-D_{2}^{\top}\end{pmatrix}=\begin{pmatrix}0&0&\tilde{a}-\alpha\\ 0&-2\alpha\tilde{b}&-\tilde{b}\alpha-1\\ \tilde{a}-\alpha&-\tilde{b}\alpha-1&-2\end{pmatrix}$.
We state that $L_{2}\preceq 0$, which is equivalent to
$0\preceq\begin{pmatrix}0&0&-\tilde{a}+\alpha\\ 0&2\alpha\tilde{b}&\tilde{b}\alpha+1\\ -\tilde{a}+\alpha&\tilde{b}\alpha+1&2\end{pmatrix}=\begin{pmatrix}0&0&0\\ 0&2&2\\ 0&2&2\end{pmatrix}$.
This is immediate. By the Positive Real Lemma, we conclude that
$\int_{0}^{T}v_{t}\tilde{y}_{t}dt=\int_{0}^{T}v_{t}(v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t})dt\geq 0,$
and the proof follows upon taking expectations. ∎
Proof of Lemma 4.2.
To proceed with the proof of Lemma 4.2, we make use of Assumption $(\boldsymbol{A_{4}})$. We fix $\epsilon,\epsilon^{\prime}>0$ small enough, along with $A\in(0,\frac{\eta}{2})$ and $h\in(0,1)$, such that
$2\sqrt{\left(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon\right)Ah}>\alpha.$ (A.4)
To apply Lemma 4.1, which requires zero initial conditions, we introduce three processes $P_{t}^{\dagger}$, $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$, whose dynamics satisfy
$d\overline{X}_{t}=Q_{t}dt+\frac{1}{N}\sum_{k=1}^{N}\sigma_{k}dW_{t}^{k}:=Q_{t}dt+\sigma dB_{t},\quad\overline{X}_{0}=\bar{x}_{0}^{N},\quad\bigg(\sigma=\frac{1}{N}\sqrt{\sum_{k=1}^{N}\sigma_{k}^{2}}\bigg),$
$d\overline{X}_{t}^{\prime}=Q_{t}dt,\quad\overline{X}_{0}^{\prime}=0,$
$d{P}_{t}^{\dagger}=\alpha(-Q_{t}-P_{t}^{\dagger})dt,\quad P_{0}^{\dagger}=0.$
$\overline{X}_{t}$, the average of $X_{t}^{i}$, contains a Brownian motion term and does not start from zero. The other two processes start at zero and have no Brownian components. From their definitions, we obtain
$\overline{X}_{t}=\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime}+\sigma B_{t},$
and a direct computation shows that
$P_{t}^{\dagger}=-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds.$ (A.5)
The dynamics and initial value of $(\overline{X}_{t}^{\prime},\ P_{t}^{\dagger})^{\top}$ coincide with those of the process in Lemma 4.1. Therefore, Lemma 4.1 implies that
$\mathbb{E}\left[\int_{0}^{T}Q_{t}(Q_{t}+P_{t}^{\dagger}+\alpha\overline{X_{t}}^{\prime})dt\right]\geq 0.$ (A.6)
Next, recall that
$J_{soc}(\boldsymbol{v})=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{T}\left(\frac{\eta}{2}\left(X_{t}^{i}-\kappa\right)^{2}+\frac{c}{2}(v_{t}^{i})^{2}+P_{t}v_{t}^{i}\right)dt+\frac{\gamma}{2}(X_{T}^{i}-\zeta)^{2}\right].$
Since (A.6) involves $P_{t}^{\dagger}$ and $\overline{X}_{t}^{\prime}$, which do not appear in $J_{soc}$, we aim to express $J_{soc}$ in a form where $\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ and part of $\frac{c}{2}v_{t}^{i}$ are replaced by terms of $P_{t}^{\dagger}$, $Q_{t}$ and $\overline{X}_{t}^{\prime}$. Our first step uses the convexity of $L$, which indicates that
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\bigg(\frac{\eta}{2}(\overline{X}_{t})^{2}-\eta\kappa\overline{X}_{t}+\frac{\eta\kappa^{2}}{2}+\Big(\frac{c}{2}-\epsilon^{\prime}\Big)Q_{t}^{2}+P_{t}Q_{t}+\epsilon^{\prime}(v_{t}^{i})^{2}\bigg)dt\right].$ (A.7)
Our second step is based on separating the control-independent part of the price. We introduce the function
$C(t)=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds,\quad\forall t\in[0,T],$
which is uniformly bounded on $[0,T]$. Using the exact form of $P_{t}$, we have
$P_{t}$ $=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds-\int_{0}^{t}e^{-\alpha(t-s)}\alpha Q_{s}ds$
$=P_{t}^{\dagger}+C(t).$ (A.8)
Our third step concerns the relationship between $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$.
$\mathbb{E}\left[(\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime})^{2}\right]$ $=\mathbb{E}\left[(\overline{X}_{t}-\sigma B_{t})^{2}\right]$
$\leq\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+\left(\frac{1}{h}-1\right)\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+\frac{h}{1-h}\mathbb{E}\left[\sigma^{2}B_{t}^{2}\right]+O(1)$
$=\frac{1}{h}\mathbb{E}\left[(\overline{X}_{t})^{2}\right]+O(1).$
This implies that
$\mathbb{E}[(\overline{X}_{t})^{2}]\geq h\mathbb{E}[(\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime})^{2}]+O(1).$ (A.9)
Substituting (A) and (A.9) into (A.7) and rearranging terms yield
$J_{soc}(\boldsymbol{v})$ $\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(\frac{\eta h}{2}\left({(\bar{x}_{0}^{N})}^{2}+2\overline{X_{t}}^{\prime}\bar{x}_{0}^{N}\right)-\eta\kappa\overline{X_{t}}^{\prime}+\left(\frac{\eta h}{2}-Ah\right)\overline{X_{t}}^{\prime 2}\right)dt\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\int_{0}^{T}\left(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X_{t}}^{\prime}+Ah\overline{X_{t}}^{\prime 2}+\left(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon\right)Q_{t}^{2}-\alpha Q_{t}\overline{X_{t}}^{\prime}\right)dt\right]$
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(C(t)Q_{t}+\epsilon Q_{t}^{2}+\epsilon^{\prime}{(v_{t}^{i})}^{2}\right)dt\right]+O(1).$ (A.10)
We have shown via Lemma 4.1 that $\mathbb{E}\left[\int_{0}^{T}(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X}_{t}^{\prime})dt\right]\geq 0$, and by (A.4) and the fundamental inequality, $(Ah\overline{X}_{t}^{\prime 2}+(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon)Q_{t}^{2}-\alpha Q_{t}\overline{X}_{t}^{\prime})$ stays non-negative. Hence we reduce (A) into
$J_{soc}(\boldsymbol{v})$ $\geq\frac{1}{N}\mathbb{E}\bigg[\sum_{i=1}^{N}\int_{0}^{T}\bigg(\frac{\eta h}{2}\Big(({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\Big)-\eta\kappa\overline{X}_{t}^{\prime}+\left(\frac{\eta h}{2}-Ah\right)\overline{X}_{t}^{\prime 2}$
$\qquad\qquad+\left(\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|\right)+\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\bigg)dt\bigg]+O(1).$
Finally, both $(\frac{\eta}{2}h-Ah)\overline{X}_{t}^{\prime 2}+\frac{\eta h}{2}\big(({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\big)-\eta\kappa\overline{X}_{t}^{\prime}$ and $\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|$ have $O(1)$ lower bounds, so
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\left(\int_{0}^{T}\left(\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\right)dt\right)\right]+O(1),$
and the proof of Lemma 4.2 is complete.
Appendix B Proofs of Auxiliary Propositions
In this section, we collect the proofs of the propositions.
Proof of section 3.
First, we need to confirm that (3.6) admits a unique classical solution on $[0,T]$. As stated before, we abbreviate $a_{i}$ and $B_{i}$ as $a$ and $B$, respectively. To prove existence, let function
$Z(a)=-\frac{\eta}{2}+\frac{2a^{2}}{c},\quad a\in\mathbb{R}.$
Note that if $\gamma\neq\sqrt{c\eta}$, (3.6) is equivalent to
$\frac{da}{Z(a)}=dt,\quad a(T)=\frac{\gamma}{2}.$
By integrating the above equation from $T$ to $t$, and observing that if $\gamma=\sqrt{c\eta}$, the solution must be a constant, we obtain that
$\left\{\begin{aligned} &a(t)=-\frac{\sqrt{c\eta}}{2}\frac{1}{\tanh\left(\sqrt{\frac{\eta}{c}}(t-T)-\operatorname{arctanh}\left(\frac{\sqrt{c\eta}}{\gamma}\right)\right)},\quad\gamma>\sqrt{c\eta},\\ &a(t)\equiv\frac{\sqrt{c\eta}}{2},\quad\gamma=\sqrt{c\eta},\\ &a(t)=\frac{\sqrt{c\eta}}{2}\operatorname{tanh}\left(\sqrt{\frac{\eta}{c}}\left(t-T\right)+\operatorname{arctanh}\left(\frac{\gamma}{\sqrt{c\eta}}\right)\right),\quad\gamma<\sqrt{c\eta}.\end{aligned}\right$.
From the analysis above, we conclude that (3.6) gives a unique solution on $[0,T]$. In the following, it is convenient to explicitly compute from (3.7)
$B(t)=e^{-A(t)}\bigg(-e^{A(T)}\gamma\zeta-\int_{t}^{T}\left(\frac{2a(s)\bar{P}(s)}{c}+\eta\kappa\right)e^{A(s)}ds\bigg),$
where $A(t)$ is a primitive function for $\frac{2a(t)}{c}$, so (3.7) gives a unique solution on $[0,T]$. The solution to (3.8) obviously exists and is unique, once we integrate both sides from $T$ to $t$. The proof is complete. ∎
Proof of section 3.
Let
$\bar{k}(t)=\begin{pmatrix}B(t)\\ \bar{x}(t)\\ \bar{P}(t)\end{pmatrix},\quad M(t)=\begin{pmatrix}\frac{2a(t)}{c}&0&\frac{2a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}\\ \frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&-\alpha+\frac{\alpha}{c}\end{pmatrix},\quad K=\begin{pmatrix}\eta\kappa\\ 0\\ \alpha\beta\end{pmatrix}$.
By Peano theorem, the initial condition gives a local solution. We need to verify that this solution extends to $t=T$ and does not blow up.
We introduce the following notation: the Frobenius norm of any matrix $L$ is denoted by $\|L\|_{\mathrm{F}}$. It is clear that $\|M(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$; we denote this bound by $M^{*}$, and denote $\|\bar{k}(t)\|_{\mathrm{F}}$ by $n(t)$. Multiply both sides of (3.9) with $2\bar{k}(t)^{\top}$, we obtain
$\frac{d}{dt}(n^{2}(t))=2(\bar{k}(t)^{\mathrm{T}}M(t)\bar{k}(t)+\bar{k}(t)^{\top}K)\leq 2n^{2}(t)M^{*}+n(t)\|K\|_{\mathrm{F}}.$
By ODE comparison theorems, $n(t)$ stays below a function $N(t)$ characterized by
$d(N^{2}(t))=2(M^{*}N^{2}(t)+\|K\|_{\mathrm{F}}N(t)),\quad N(0)=n(0)+1.$
It directly follows that $N(t)$ has a uniform bound on $t\in[0,T]$, so $\bar{k}(t)$ stays finite on $[0,T]$. Therefore, the existence of the solution holds. Moreover, $M(t)$ is Lipschitz continuous w.r.t $t\in[0,T]$, so uniqueness naturally follows from the Picard-Lindelöf theorem. The proof is complete.
Proof of section 3.
We adopt the notation established in the proofs of section 3. Deducing in the same way as in the proof of section 3, we obtain the existence property. Recalling that $M(t)$ is Lipschitz continuous, we obtain the uniqueness property. ∎
Proof of section 3.
We adopt the notation established previously in section 3. By the definition of $\phi_{b_{0}}$, as shown in section 3 , we discover that for any $b_{0}\neq 0$, $\frac{\phi_{b_{0}}-\phi_{0}}{b_{0}}$ solves (3.11). Consequently, from the uniqueness property provided by section 3, we have
$\phi_{b_{0}}(t)-\phi_{0}(t)=b_{0}B_{1}(t).$
Set $t=T$. From $B_{1}(T)\neq 0$, we conclude that there exists a unique $b_{0}\in\mathbb{R}$ that satisfies $\phi_{b_{0}}(T)=-\zeta\gamma$. Thus, there exists only one solution to (3.9) and (3.10). The proof is complete. ∎
Proof of section 4.
Set
$G(t)=\begin{pmatrix}-\alpha+\frac{\alpha}{c}&\frac{2\alpha a(t)}{c}&\frac{\alpha}{c}&\frac{\alpha^{2}}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&-\frac{\alpha}{c}\\ \frac{2a(t)}{c}&0&\frac{2a(t)}{c}&\frac{2\alpha a(t)}{c}\\ -\frac{1}{c}&-\frac{2a(t)}{c}&-\frac{1}{c}&\alpha-\frac{\alpha}{c}\end{pmatrix}$.
To prove existence, we observe that $\|G(t)\|_{\mathrm{F}}$ has a uniform bound on $[0,T]$. Consequently, by working exactly as in section 3, we obtain the uniqueness property. Note that $G(t)$ is continuously differentiable, so the solution of the system is unique, from the Picard-Lindelöf theorem. ∎
Proof of section 4.
The existence holds by the same way as in section 3. Recalling from section 4 that $G(t)$ is continuously differentiable everywhere, we conclude that the solution is unique, by the Picard-Lindelöf theorem. ∎
Proof of section 4.
Observe that , for $\forall b_{0},\ l_{0}\in\mathbb{R}$, $b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$ solves (4.11), and has boundary values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$. Then, uniqueness property in section 4 yields that $\phi_{b_{0},l_{0}}^{*}=b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$. Therefore, by setting $t=T$, we get $\phi_{b_{0},l_{0}}^{*}(T)=b_{0}\phi_{1}^{*}(T)+l_{0}\phi_{2}^{*}(T)+\phi_{0,0}^{*}(T)$. By Assumption $(\boldsymbol{A_{3}})$, there exists a unique pair of constants $b_{0},\ l_{0}\in\mathbb{R}$ such that the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a solution that satisfies (4.11) and (4.12). Thus, the solution for (4.11) and (4.12) exists and is unique. The proof is complete. ∎